### v18

# Simulation of Polynomial-Time Nondeterministic Turing Machines via Universal Nondeterministic Turing Machines

Abstract.
We prove in this paper that there is a language $L_{s}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$ machines. Then we further show that $L_{s}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are simulation and the novel new techniques developed in the author’s recent work. Further, if there exists some oracle $A$ such that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery lies behind it and show that if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable, thus showing that the simulation techniques are not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. Our main result has profound implications, such as $\mathcal{P}\neq\mathcal{NP}$ and a lower bounds result for Frege proof systems (i.e., no Frege proof systems can be polynomially bounded), etc.
keywords:
$\mathcal{NP}$
keywords:
Polynomial-time nondeterministic Turing machine
keywords:
${\rm co}\mathcal{NP}$
keywords:
Simulation
keywords:
Universal nondeterministic Turing machine
keywords:
Frege system
1. Introduction
Let $\mathcal{NP}$ be the set of decision problems solvable in polynomial time by a nondeterministic Turing machine, i.e.,
[…]
$\qquad\,\,\text{that accepts the language $L$}\big{\}}.$
Let ${\rm co}\mathcal{NP}$ be the complement of $\mathcal{NP}$, i.e.,
${\rm co}\mathcal{NP}=\left\{\overline{L}\,:\,L\in\mathcal{NP}\right\}$
where $\overline{L}$ is the set-theoretic complement of $L$ (i.e., assume $L\subseteq\{0,1\}^{*}$, then $\overline{L}=\{0,1\}^{*}\setminus L$). The outstanding and long-standing open problem $\mathcal{NP}\overset{?}{=}{\rm co}\mathcal{NP}$ asks whether $\mathcal{NP}$ equals ${\rm co}\mathcal{NP}$ or not (i.e., the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem) and is a central topic in computational complexity theory, whose implications are profound for computer science, mathematics, and related fields.
To see the profound implications in various aspects of this conjecture, let us dive into two cases: one is $\mathcal{NP}={\rm co}\mathcal{NP}$, and another is $\mathcal{NP}\neq{\rm co}\mathcal{NP}$.
If $\mathcal{NP}={\rm co}\mathcal{NP}$, then for every problem where “yes” answers are efficiently verifiable, “no” answers would also be efficiently verifiable. This would be surprising because many problems, such as Satisfiability [Coo71, Lev73] or Graph non-isomorphism [GJ79], seem to lack obvious short certificates for “no” instances. It is also widely believed that $\mathcal{NP}={\rm co}\mathcal{NP}$ would imply significant consequences for the Polynomial hierarchy ($PH$) [Sto77, A6], a layered structure of complexity classes built on $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$. Specifically, if $\mathcal{NP}={\rm co}\mathcal{NP}$, then $PH$ would collapse to $\mathcal{NP}$, i.e., $PH=\Sigma_{1}^{P}=\prod_{1}^{P}=\mathcal{NP}$; see e.g. Theorem 5.4 in [AB09]. This contradicts the widely believed assumption that $PH$ is infinitely extended. Note that this conjecture is not only one of the holy grail problems of theoretical computer science but also the core key to understanding the essence of computing.
Furthermore, many cryptographic systems rely on problems such as Integer factorization [RSA78] and Discrete logarithm [BG04] being hard to solve but easy to verify. If $\mathcal{NP}={\rm co}\mathcal{NP}$, then the “no” instances might also become easy to verify, potentially weakening certain cryptographic assumptions. If $\mathcal{NP}={\rm co}\mathcal{NP}$, then optimization problems tied to $\mathcal{NP}$ (such as finding the shortest path in Traveling Salesman Problem (TSP) [Law85]) would have complements (e.g., proving no shorter path exists) that are equally tractable, which could revolutionize industries like logistics, where verifying optimality becomes as easy as finding solutions.
However, if $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, it reinforces the belief that there is an asymmetry in computing: problems with efficiently verifiable “yes” instances do not necessarily have efficiently verifiable “no” instances. In particular, it would imply that $\mathcal{P}$ and $\mathcal{NP}$ differ. Furthermore, if $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, the status quo in cryptography remains. The asymmetry ensures that problems can be hard to solve yet easy to verify in one direction, supporting the foundations of modern security protocols. Simultaneously, verifying the absence of a solution remains potentially intractable, meaning that we might efficiently find solutions but struggle to prove they are optimal, a common challenge in real-world applications.
Thus far, although the exact relationship between the complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ is unknown and the complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ are widely believed to be different, Figure 1 below illustrates the possibility of the relationship between the complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ that most complexity theorists believed.
Figure 1. The most believed possibility between $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$
Our motivation in this paper is that, in the author’s recent work [Lin21], we left this outstanding and long-standing open conjecture $\mathcal{NP}\overset{?}{=}{\rm co}\mathcal{NP}$ untouched on. In the present paper, we try to resolve this important open conjecture.
In addition, the $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ conjecture is closely related to proof complexity. The study of proof systems was initially motivated by the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem. Cook and Reckhow [CR79] define a proof system for a set $L$ as a polynomial-time computable function $f$ whose range is $L$. They show that $\mathcal{NP}$ if and only if there exists a proof system with polynomial-size proofs for the set of propositional tautologies.
As our second goal, we will consider the proof systems that are familiar from textbook presentations of logic, such as the Frege systems. It can be said that the Frege proof system is a “textbook-style” propositional proof system with Modus Ponens as its only rule of inference. Interestingly, there is also an important open problem asking to show a superpolynomial lower bound on the length of proofs for a Frege system or to prove that it is polynomially bounded. Namely, the following
Prove a superpolynomial lower bound on the length of proofs for a Frege system (or prove that it is polynomially bounded).
Indeed, the above problem was listed as the first open problem in [Pud08]. However, despite decades of effort, very little progress has been made on this long-standing open problem. Note that, previously, although $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ is considered to be very likely true, researchers are not able to prove that some very basic proof systems are not polynomially bounded; see e.g. [Pud08]. Our second goal is to resolve this long-standing problem.
1.1. Our Contributions
In this paper, we explore and settle the aforementioned open problems. Our first main goal in this paper is to prove the following most important theorem:
Theorem 1.1.
There is a language $L_{s}$ accepted by a nondeterministic Turing machine but by no ${\rm co}\mathcal{NP}$ machines, i.e., $L_{s}\notin{\rm co}\mathcal{NP}$. Further, it can be proved that $L_{s}\in\mathcal{NP}$. That is,
$\mathcal{NP}\neq{\rm co}\mathcal{NP}.$
The notion of ${\rm co}\mathcal{NP}$ machines appearing in Theorem 1.1 is defined in Section 2.
Note that we prove the first half of Theorem 1.1 by the ordinary simulation techniques and the relationship between languages $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and the complement $\overline{L}(M)$ of $L(M)$. Furthermore, by the result of [BGS75], since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, and just as the result that $\mathcal{P}^{A}=\mathcal{NP}^{A}$ [BGS75] led to a strong belief that problems with contradictory relativization are very hard to solve and are not amenable to current proof techniques, i.e., the solutions of such problems are beyond the current state of mathematics (see e.g. [HCCRR93, Hop84]), the conclusion $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ also suggests that the problem of separating $\mathcal{NP}$ from ${\rm co}\mathcal{NP}$ is beyond the current state of mathematics. Therefore, the next result, i.e., the following theorem, is to break the so-called Relativization Barrier:
Theorem 1.2.
Under some rational assumptions (see Section 6 below), and if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, then the set of all ${\rm co}\mathcal{NP}^{A}$ machines (defined in Section 6) is not enumerable. Thereby, ordinary simulation techniques will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
Meanwhile, there is another case, i.e., the case that $\mathcal{P}^{A}=\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$. In this scenario, we remark that Theorem 6.1 in Section 6 perfectly expounds the reason why we are unable to apply simulation techniques to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$.
As mentioned earlier, Theorem 1.1 has profound implications, one of which is the following  [^1] In addition, if the complexity class $\mathcal{BQP}$ (bounded-error quantum polynomial-time; see page 213 of [AB09]) has the similar property (i.e., $\mathcal{BQP}=\rm{co}\mathcal{BQP}$), then obviously we have another important relation $\mathcal{BQP}\neq\mathcal{NP}$.
Corollary 1.3.
$\mathcal{P}\neq\mathcal{NP}$.
Thus, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again.
Let $\mathcal{NEXP}$ and ${\rm co}\mathcal{NEXP}$ be the complexity classes defined by
$\mathcal{NEXP}\overset{\rm def}{=}\bigcup_{k\in\mathbb{N}_{1}}{\rm NTIME}[2^{n^{k}}]$
and
${\rm co}\mathcal{NEXP}\overset{\rm def}{=}\bigcup_{k\in\mathbb{N}_{1}}{\rm coNTIME}[2^{n^{k}}],$
respectively (the complexity classes ${\rm NTIME}$ and ${\rm coNTIME}$ are defined below, see Section 2). Then, besides the above obvious corollary, we also have the following consequence of Theorem 1.1, whose proof can be similar to that of Theorem 1.1:
Corollary 1.4.
$\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$.
Since $\mathcal{EXP}={\rm co}\mathcal{EXP}$ (where $\mathcal{EXP}$ is the class of languages accepted by deterministic Turing machines in time $2^{n^{O(1)}}$; see Section 2 for big $O$ notation), this, together with Corollary 1.4, follows that $\mathcal{EXP}\neq\mathcal{NEXP}$. Furthermore, the complexity class $\mathcal{BPP}$ (bounded-error probabilistic polynomial-time) aims to capture the set of decision problems efficiently solvable by polynomial-time probabilistic Turing machines (see e.g. [AB09]), and a central open problem in probabilistic complexity theory is the relation between $\mathcal{BPP}$ and $\mathcal{NEXP}$. However, currently researchers only know that $\mathcal{BPP}$ is sandwiched between $\mathcal{P}$ and $\mathcal{EXP}$ (i.e., $\mathcal{P}\subseteq\mathcal{BPP}\subseteq\mathcal{EXP}$) but are even unable to show that $\mathcal{BPP}$ is a proper subset of $\mathcal{NEXP}$; see e.g. page 126 of [AB09]. With the consequence of Corollary 1.4 (i.e., $\mathcal{EXP}\neq\mathcal{NEXP}$) at hand, it immediately follows that
Corollary 1.5.
$\mathcal{BPP}\neq\mathcal{NEXP}$.
It is interesting that the complexity class $\mathcal{NP}$ has a rich structure if $\mathcal{P}$ and $\mathcal{NP}$ differ. Specifically, in [Lad75], Lander constructed a language that is $\mathcal{NP}$-intermediate under the assumption that $\mathcal{P}\neq\mathcal{NP}$. In fact, since by Theorem 1.1, we know that $\mathcal{P}\neq{\rm co}\mathcal{NP}$ and symmetrically also have the following interesting outcome saying that the complexity class ${\rm co}\mathcal{NP}$ has the intermediate languages. To see so, let $L\in\mathcal{NP}$ be the $\mathcal{NP}$-intermediate language constructed in [Lad75]; then $\overline{L}=\{0,1\}^{*}-L$ is a ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$. In other words, we will prove in detail the following important result:
Theorem 1.6.
There are ${\rm co}\mathcal{NP}$-intermediate languages, i.e., there exists language $L\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete.
By the result of [CR74], there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement. Then, by Theorem 1.1, we clearly have the following:
Corollary 1.7.
There exists no super proof system.
Moreover, we will settle the open problem listed as the first in [Pud08]; see the introduction in Section 1. Namely, we show the following theorem with respect to the aforementioned problem:
Theorem 1.8.
There exists no polynomial $p(n)$ such that for all $\psi\in TAUT$, there is a Frege proof of $\psi$ of length at most $p(|\psi|)$. In other words, no Frege proof systems of $\psi\in TAUT$ can be polynomially bounded.
1.2. Background and Prior Work
In an epoch-making $36$-page paper [Tur37] published in 1936, Turing opened the door to the computer science. Turing’s contributions, on the one hand, were highly influential in the development of theoretical computer science that he is widely regarded as the father of theoretical computer science [A7]. But on the other hand, although Turing’s work initiated the study of theoretical computer science, he was not concerned with the efficiency of his machines, which is the main topic in computational complexity theory. In fact, Turing’s concern [Tur37] was whether they can simulate arbitrary algorithms given sufficient time (see e.g. [Coo00]).
Why are some problems easy for computers to solve, while others seem fundamentally intractable? This question lies at the heart of computational complexity theory. So what is computational complexity theory? Computational complexity theory is a central subfield of theoretical computer science and mathematics, which mainly concerns the efficiency of Turing machines (i.e., algorithms) or the intrinsic complexity of computational tasks, i.e., focuses on classifying computational problems according to their resource usage and relating these classes to each other (see e.g. [A1]). In other words, computational complexity theory specifically deals with fundamental questions such as what is feasible computation and what can and cannot be computed with a reasonable amount of computational resources in terms of time or space (memory). What’s exciting is that there are many remarkable open problems in the area of computational complexity theory, such as the famous $\mathcal{P}$ versus $\mathcal{NP}$ problem and the important unknown relation between the complexity class $\mathcal{NP}$ and the complexity class ${\rm co}\mathcal{NP}$, which deeply attract many researchers to conquer them. However, despite decades of effort, very little progress has been made on these important problems. Furthermore, as pointed out by Wigderson [Wig07], to understand the power and limits of efficient computation has led to the development of computational complexity theory, and this discipline in general, and the aforementioned famous open problems in particular, have gained prominence within the mathematics community in the past decades.
As introduced in Wikipedia [A1], in the field of computational complexity theory, a problem is regarded as inherently difficult if its solution requires significant resources such as time and space, whatever the algorithm used. The theory formalizes this intuition by introducing mathematical models of computing to study these problems and quantifying their computational complexity, i.e., the amount of resources needed to solve them. However, it is worth noting that other measures of complexity are also used, such as the amount of communication, which appeared in communication complexity, the number of gates in a circuit, which appeared in circuit complexity, and so forth (see e.g. [A1]). Going further, one of the roles of computational complexity theory is to determine the practical limits on what computers (the computing models) can and cannot do.
In history, the fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. Specifically, $\mathcal{NP}$ is the set of decision problems for which the problem instances, where the answer is “yes,” have proofs verifiable in polynomial time by some deterministic Turing machine, or alternatively the set of problems that can be solved in polynomial time by a nondeterministic Turing machine (see e.g. [A2]). The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that the Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial time. At the same time, it is also worth noting that this famous and seminal result opened the door to research into the renowned rich theory of $\mathcal{NP}$-completeness [Kar72].
On the other hand, the complexity class ${\rm co}\mathcal{NP}$ is formally defined as follows: For a complexity class $\mathcal{C}$, its complement is denoted by ${\rm co}\mathcal{C}$ (see e.g. [Pap94]), i.e.,
${\rm co}\mathcal{C}=\{\overline{L}:L\in\mathcal{C}\},$
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$ (i.e., $\overline{L}=\Sigma^{*}\setminus L$ with the assumption that the language $L$ is over the alphabet $\Sigma$). That is to say, ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa. In other words, for instance, according to Wikipedia [A3], ${\rm co}\mathcal{NP}$ is the set of decision problems where there exists a polynomial $p(n)$ and a polynomial-time bounded Turing machine $M$ such that for every instance $x$, $x$ is a no-instance if and only if: for some possible certificate $c$ of length bounded by $p(n)$, the Turing machine $M$ accepts the pair $(x,c)$. To the best of our knowledge, the complexity class ${\rm co}\mathcal{NP}$ was introduced for the first time by Meyer and Stockmeyer [MS72] with the name “$\prod_{1}^{P}$,” and Stockmeyer wrote a full paper on the polynomial hierarchy, which also uses the notation ${\rm co}\mathcal{NP}$; see e.g. [Sto77].
The $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is deeply connected to the field of proof complexity. Here are some historical notes: In 1979, Cook and Reckhow [CR79] introduced a general definition of a propositional proof system and related it to mainstream complexity theory by pointing out that such a system exists in which all tautologies have polynomial-length proofs if and only if the two complexity classes $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ coincide; see e.g. [CN10]. We refer the reader to the reference [Coo00] for the importance of this research field and to the reference [Kra95] for the motivation of the development of this rich theory. Apart from those, Chapter $10$ of [Pap94] also contains the introductions of the importance of the problem
$\mathcal{NP}\overset{?}{=}{\rm co}\mathcal{NP}.$
In the next few paragraphs, let us recall some of the interesting background, the current status, the history, and the main goals of the field of proof complexity. On the one hand, as we all know well, proof theory is a major branch of mathematical logic and theoretical computer science with which proofs are treated as formal mathematical objects, facilitating their analysis by mathematical techniques (see e.g. [A4]). Some of the major areas of proof theory include structural proof theory, ordinal analysis, automated theorem proving, and proof complexity, and so forth. It should be pointed out that much research also focuses on applications in computer science, linguistics, and philosophy, from which we see that the impact of proof theory is very profound. Moreover, as pointed out in [Raz04] by Razborov, proof and computations are among the most fundamental concepts relevant to virtually any human intellectual activity. Both have been central to the development of mathematics for thousands of years. The effort to study these concepts themselves in a rigorous, metamathematical way initiated in the $20$th century and led to the flourishing of mathematical logic and derived disciplines.
On the other hand, according to the points of view given in [Rec76], logicians have proposed a great number of systems for proving theorems, which give certain rules for constructing proofs and for associating a theorem (formula) with each proof. More importantly, these rules are much simpler to understand than the theorems. Thus, a proof gives a constructive way of understanding that a theorem is true. Specifically, a proof system is sound if every theorem is true, and it is complete if every true statement (from a certain class) is a theorem (i.e., has a proof); see e.g., [Rec76].
Indeed, the relationship between proof theory and proof complexity is inclusion. Within the area of proof theory, proof complexity is the subfield aiming to understand and analyze the computational resources that are required to prove or refute statements (see e.g., [A5]). Research in proof complexity is predominantly concerned with proving proof-length lower and upper bounds in various propositional proof systems. For example, among the major challenges of proof complexity is showing that the Frege system, the usual propositional calculus, does not admit polynomial-size proofs of all tautologies; see e.g. [A5]. Here the size of the proof is simply the number of symbols in it, and a proof is said to be of polynomial size if it is polynomial in the size of the tautology it proves. Moreover, contemporary proof complexity research draws ideas and methods from many areas in computational complexity, algorithms, and mathematics. Since many important algorithms and algorithmic techniques can be cast as proof search algorithms for certain proof systems, proving lower bounds on proof sizes in these systems implies run-time lower bounds on the corresponding algorithms; see e.g. [A5].
Likewise, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see e.g. the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g. [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see e.g. [Raz03] for interesting directions in propositional proof complexity.
Historically, systematic study of proof complexity began with the seminal work of Cook and Reckhow [CR79], who provided the basic definition of a propositional proof system from the perspective of computational complexity. In particular, as pointed out by the authors of [FSTW21], the work [CR79] relates the goal of propositional proof complexity to fundamental hardness questions in computational complexity such as the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem: establishing super-polynomial lower bounds for every propositional proof system would separate $\mathcal{NP}$ from ${\rm co}\mathcal{NP}$.
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem for this moment. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see e.g. the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in our, and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g. [Kra19].
Particularly, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any super proof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems, see e.g. the excellent survey on proof complexity [Bus12].
1.3. Our Approach
The aforementioned “Cook’s program” is among the most well-known programs to resolve the long-standing open conjecture $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems.
As a novel idea and attempt, we will apply simulation techniques. To the best of our knowledge, if we want to create a language not in a specific complexity class $\mathcal{C}$, the most obvious way is to apply the diagonalization technique or the lazy-diagonalization technique, because the simulation and diagonalization technique [Tur37, HS65] is a standard method (currently the only way) to prove lower bounds on uniform computing models (i.e., the Turing machine model). But, we find that it is hard to apply standard diagonalization or lazy-diagonalization techniques to produce a language that not in ${\rm co}\mathcal{NP}$. Instead, by the relation between the language $L(M)$ accepted by polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by ${\rm co}\mathcal{NP}$ machine $M$, we will apply only simulation techniques. So, we first try to show that all polynomial-time nondeterministic Turing machines are enumerable. Then, we use a universal nondeterministic Turing machine to simulate all polynomial-time nondeterministic Turing machines, creating a language that is not in ${\rm co}\mathcal{NP}$. To stress it is “novel,” it is because this method is the first appearing in this paper, and we have never seen such an attempt appeared in previous works or such an idea in the relevant standard textbooks, such as [AHU74, AB09, Sip13, Pap94].
After doing so, we further prove that such a language is actually in $\mathcal{NP}$ by using the novel method first appearing in the author’s recent work [Lin21], thus obtaining the desired result $\mathcal{NP}\neq{\rm co}\mathcal{NP}$.
1.4. Organization
The remainder of this paper is organized as follows: For the convenience of the reader, we will review some notions in the next section closely associated with our discussions and fix some notation we will use in the following context, with some useful technical lemmas also presented in the next section. In Section 3, we prove that all polynomial-time nondeterministic Turing machines are enumerable and all ${\rm co}\mathcal{NP}$-machines are enumerable, thus showing that there exists an enumeration of all ${\rm co}\mathcal{NP}$-machines. Section 4 contains the definition of our universal nondeterministic Turing machine, which accepts a language $L_{s}$ not in ${\rm co}\mathcal{NP}$, showing the desired lower bounds. And the Section 5 is devoted to showing that the language $L_{s}$ is in fact in $\mathcal{NP}$, proving the upper bounds required, in which we finish the proof of Theorem 1.1 at our own convenience. In Section 6 we will show Theorem 1.2 to break the so-called Relativization Barrier. In Section 7, we show that there are ${\rm co}\mathcal{NP}$-intermediate languages in the complexity class ${\rm co}\mathcal{NP}$. In Section 8, we prove the Theorem 1.8 which says that no Frege proof systems are polynomially bounded, thus answering an important open question in the area of proof complexity. Finally, we list some open problems for future research in the last section.
2. Preliminaries
In this section, we introduce some notions and notation that will be used in what follows.
Let $\mathbb{N}_{1}=\{1,2,3,\cdots\}$ where $+\infty\notin\mathbb{N}_{1}$, i.e., the set of all positive integers. We also denote by $\mathbb{N}$ the set of natural numbers, i.e.,
=$\mathbb{N}\overset{\rm def}{=}\mathbb{N}_{1}\cup\{0\}.$
Let $\Sigma$ be an alphabet, for finite words $w,v\in\Sigma^{*}$, the concatenation of $w$ and $v$, denoted by $w\circ v$, is $wv$. For example, suppose that $\Sigma=\{0,1\}$, and $w=100001$, $v=0111110$, then
$w\circ v\overset{\rm def}{=}$ $wv$
$=$ 10000101111101000010111110 10000101111101000010111110
The length of a finite word $w$, denoted as $|w|$, is defined to be the number of symbols in it. It is clear that for finite words $w$ and $v$,
$|wv|=|w|+|v|.$
The big $O$ notation indicates the order of growth of some quantity as a function of $n$ or the limiting behavior of a function. For example, that $S(n)$ is big $O$ of $f(n)$, i.e.,
$S(n)=O(f(n))$
means that there exist a positive integer $N_{0}$ and a positive constant $M$ such that
$S(n)\leq M\times f(n)$
for all $n>N_{0}$.
The little $o$ notation also indicates the order of growth of some quantity as a function of $n$ or the limiting behavior of a function but with different meaning. Specifically, that $T(n)$ is little $o$ of $t(n)$, i.e.,
$T(n)=o(t(n))$
means that for any constant $c>0$, there exists a positive integer $N_{0}>0$ such that
$T(n)<c\times t(n)$
for all $n>N_{0}$.
Throughout this paper, the computational modes used are nondeterministic Turing machines (or their variants such as nondeterministic Turing machines with oracle). We follow the standard definition of a nondeterministic Turing machine given in the standard textbook [AHU74]. Let us first introduce the precise definition of a nondeterministic Turing machine as follows:
Definition 2.1 ($k$-tape nondeterministic Turing machine, [AHU74]).
A $k$-tape nondeterministic Turing machine (shortly, NTM) $M$ is a seven-tuple $(Q,T,I,\delta,\mathbbm{b},q_{0},q_{f})$ where:
  1. (1)
    $Q$ is the set of states.
  2. (2)
    $T$ is the set of tape symbols.
  3. (3)
    $I$ is the set of input symbols; $I\subseteq T$.
  4. (4)
    $\mathbbm{b}\in T-I$, is the blank.
  5. (5)
    $q_{0}$ is the initial state.
  6. (6)
    $q_{f}$ is the final (or accepting) state.
  7. (7)
    $\delta$ is the next-move function, or a mapping from $Q\times T^{k}$ to subsets of
    $Q\times(T\times\{L,R,S\})^{k}.$
    Suppose
    $\delta(q,a_{1},a_{2},\cdots,a_{k})=\{(q_{1},(a^{1}_{1},d^{1}_{1}),(a^{1}_{2},d^{1}_{2}),\cdots,(a^{1}_{k},d^{1}_{k})),\cdots,(q_{n},(a^{n}_{1},d^{n}_{1}),(a^{n}_{2},d^{n}_{2}),\cdots,(a^{n}_{k},d^{n}_{k}))\}$
    and the nondeterministic Turing machine is in state $q$ with the $i$th tape head scanning tape symbol $a_{i}$ for $1\leq i\leq k$. Then in one move the nondeterministic Turing machine enters state $q_{j}$, changes symbol $a_{i}$ to $a^{j}_{i}$, and moves the $i$th tape head in the direction $d^{j}_{i}$ for $1\leq i\leq k$ and $1\leq j\leq n$.
Let $M$ be a nondeterministic Turing machine, and $w$ be an input. Then $M(w)$ represents that $M$ is on input $w$.
A nondeterministic Turing machine $M$ works in time $t(n)$ (or of time complexity $t(n)$), if for any input $w\in I^{*}$ where $I$ is the input alphabet of $M$, $M(w)$ will halt within $t(|w|)$ steps.
Now, it is time for us to introduce the concept of a polynomial-time nondeterministic Turing machine as follows:
Definition 2.2 (cf. polynomial-time deterministic Turing machines in [Coo00]).
Formally, a polynomial-time nondeterministic Turing machine is a nondeterministic Turing machine such that there exists $k\in\mathbb{N}_{1}$, for all input $w$ of length $|w|$ where $|w|\in\mathbb{N}$, $M(w)$ will halt within $t(|w|)=|w|^{k}+k$ steps. The language accepted by polynomial-time nondeterministic Turing machine $M$ is denoted by $L(M)$.
By default, a word $w$ is accepted by a polynomial-time (say, $t(n)$ time-bounded) nondeterministic Turing machine $M$ if there exists at least one computation path $\gamma$ putting the machine into an accepting state (i.e., stopping in the “accepting” state) on input $w$ (in such a situation, we call $M$ accepts $w$). Specifically, $M$ accepts the language $L(M)$, then $w\in L(M)$ if and only if $M$ accepts $w$. In other words, there exists at least one accepting path for $M$ on input $w$. This is the “there exists at least one accepting path” condition for nondeterministic Turing machines, based on which the complexity class ${\rm NTIME}[t(n)]$ is defined:
Definition 2.3 (“there exists at least one accepting path” condition).
The set of languages decided by nondeterministic Turing machines within time $t(n)$ by the condition of “there exists at least one accepting path” is denoted by ${\rm NTIME}[t(n)]$. Under this condition, a $t(n)$ time-bounded nondeterministic Turing machine accepts a string $w$ if and only if there exists at least one accepting path on input $w$. Thus,
$\mathcal{NP}=\bigcup_{k\in\mathbb{N}_{1}}{\rm NTIME}[n^{k}].$
Apart from the “there exists at least one accepting path” condition for defining the complexity class ${\rm NTIME}[t(n)]$, there is another condition (i.e., there exists no computation path $\gamma$ putting the machine into an accepting state), which is the dual of the “there exists at least one accepting path” condition. In what follows, we call it the “there exists no accepting path” condition.
Let $M$ be a polynomial-time (say, $t(n)$ time-bounded) nondeterministic Turing machine over the input alphabet $\{0,1\}$. Then, under the “there exists no accepting path” condition, a word $w$ is rejected by $M$ if there exists no computation path $\gamma$ putting the machine into an accepting state on input $w$ (in such a situation, we call $M$ rejects $w$). Clearly, $M$ will define the language $\overline{L}(M)$, which is the complement of $L(M)$ (i.e., $\overline{L}(M)=\{0,1\}^{*}\setminus L(M)$). Specifically, $w\in\overline{L}(M)$ if and only if $M$ rejects $w$.
Since ${\rm co}\mathcal{NP}=\{\overline{L}\,:\,L\in\mathcal{NP}\}$, a polynomial-time nondeterministic Turing machine $M$ accepts the language $L$ under the condition of “there exists at least one accepting path,” the same machine $M$ will simultaneously define the language $\overline{L}$ under the condition of “there exists no accepting path.” Thus, we can conveniently define the notion of a ${\rm co}\mathcal{NP}$ machine as follows:
Definition 2.4.
A ${\rm co}\mathcal{NP}$ machine $M$ is exactly a polynomial-time nondeterministic Turing machine $M$ working under the “there exists no accepting path” condition. We define the language accepted by ${\rm co}\mathcal{NP}$ machine $M$ to be $\overline{L}(M)$.222 In short, there is an interesting relationship: ${\rm co}\mathcal{NP}$ machine $M$ accepts the language $\overline{L}(M)$ but polynomial-time nondeterministic Turing machine $M$ rejects $\overline{L}(M)$; polynomial-time nondeterministic Turing machine $M$ accepts the language $L(M)$ but ${\rm co}\mathcal{NP}$ machine $M$ rejects $L(M)$. Namely, for any input $w$, $w\in\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ rejects $w$.
Specifically, assume that the $t(n)$ time-bounded ${\rm co}\mathcal{NP}$ machine $M$ (over the input alphabet $\{0,1\}$) accepts the language $\overline{L}(M)$. Clearly, $w\notin\overline{L}(M)$ if and only if polynomial-time nondeterministic Turing machine $M$ accepts $w$.
For a ${\rm co}\mathcal{NP}$ machine $M$, we define the characteristic function
$\chi_{\overline{L}(M)}:\{0,1\}^{*}\rightarrow\{0,1\}$
as follows:
$\chi_{\overline{L}(M)}(w)=\left\{\begin{array}[]{ll}1,&\hbox{$w\in\overline{L}(M)$;}\\ 0,&\hbox{$w\notin\overline{L}(M)$.}\end{array}\right.$
Let $M$ be a polynomial-time nondeterministic Turing machine. Since $M$ can be regarded as both a polynomial-time nondeterministic Turing machine working under the “there exists at least one accepting path” condition and a ${\rm co}\mathcal{NP}$ machine working under the “there exists no accepting path” condition, we strictly distinguish, in the following context, between these two situations in which the use of symbols. Namely, we use the notation
$M(w)=1\quad\text{(respectively, $M(w)=0$)}$
to mean that polynomial-time nondeterministic Turing machine $M$ accepts (respectively, rejects) the input $w$; we use the notation
$\chi_{\overline{L}(M)}(w)=1\quad\text{ (respectively, $\chi_{\overline{L}(M)}(w)=0$)}$
to mean that ${\rm co}\mathcal{NP}$ machine $M$ accepts (respectively, rejects) the input $w$.
It is clear that for all $w\in\{0,1\}^{*}$, $w\in\overline{L}(M)$ if and only if
$M(w)=0,\quad\text{(i.e., polynomial-time nondeterministic Turing machine $M$ rejects $w$)}$
and $w\notin\overline{L}(M)$ if and only if
$M(w)=1,\quad\text{(i.e., polynomial-time nondeterministic Turing machine $M$ accepts $w$).}$.
As mentioned earlier, the “there exists no accepting path” condition is the dual of the “there exists at least one accepting path” condition, based on which the complexity class ${\rm coNTIME}[t(n)]$ is defined.
Definition 2.5 (“there exists no accepting path” condition).
The set of languages decided by nondeterministic Turing machines within time $t(n)$ by the condition of “there exists no accepting path” is denoted by ${\rm coNTIME}[t(n)]$. Thus,
${\rm co}\mathcal{NP}=\bigcup_{k\in\mathbb{N}_{1}}{\rm coNTIME}[n^{k}].$
In fact, we can define the complexity class ${\rm co}\mathcal{NP}$ in a equivalent way using polynomial-time deterministic Turing machines as verifiers and “for all” witnesses, given as follows:
Definition 2.6 ([AB09], Definition 2.20).
For every $L\subseteq\{0,1\}^{*}$, we say that $L\in{\rm co}\mathcal{NP}$ if there exists a polynomial $p(n)$ and a deterministic polynomial-time Turing machine $M$ such that for every $x\in\{0,1\}^{*}$,
$x\in L\Leftrightarrow\forall u\in\{0,1\}^{p(|x|)},\quad M(x,u)=1.$
In regard to the relation between the time complexity of $k$-tape nondeterministic Turing machines and that of single-tape nondeterministic Turing machines, we quote the following useful lemma, extracted from the standard textbook [AHU74] (see Lemma 10.1 in [AHU74]), which plays important roles in the following context:
Lemma 2.7 (Lemma 10.1 in [AHU74]).
If $L$ is accepted by a $k$-tape nondeterministic $T(n)$ time-bounded Turing machine, then $L$ is accepted by a single-tape nondeterministic $O(T^{2}(n))$ time-bounded Turing machine.
Since ${\rm co}\mathcal{NP}$ machines are also polynomial-time nondeterministic Turing machines working under the “there exists no accepting path” condition, we can similarly prove the following technical lemma:
Lemma 2.8.
If $L$ is accepted by a $k$-tape $T(n)$ time-bounded ${\rm co}\mathcal{NP}$ machine, then $L$ is accepted by a single-tape $O(T^{2}(n))$ time-bounded ${\rm co}\mathcal{NP}$ machine.
Proof 2.9.
The proof is similar to that of Lemma 2.7 and see [AHU74], so the detail is omitted.
The following theorem about efficient simulation for universal nondeterministic Turing machines is useful in proving the main result in Section 4.
Lemma 2.10 ([AB09]).
There exists a Turing machine $U$ such that for every $x,\alpha\in\{0,1\}^{*}$, $U(x,\alpha)=M_{\alpha}(x)$, where $M_{\alpha}$ denotes the Turing machine represented by $\alpha$. Moreover, if $M_{\alpha}$ halts on input $x$ within $T(|x|)$ steps, then $U(x,\alpha)$ halts within $cT(|x|)\log T(|x|)$  [^3] In this paper, $\log n$ stands for $\log_{2}n$. steps, where $c$ is a constant independent of $|x|$ and depending only on $M_{\alpha}$’s alphabet size, number of tapes, and number of states.
Other background information and notions will be given along the way in proving our main results stated in Section 1.
3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
In this section, our main goal is to establish an important theorem that all polynomial-time nondeterministic Turing machines and all ${\rm co}\mathcal{NP}$ machines are in fact enumerable, which is the prerequisite for applying the simulation technique.
Following Definition 2.2, a polynomial-time nondeterministic Turing machine, say a $n^{k}+k$ time-bounded nondeterministic Turing machine, can be represented by the tuple of $(M,k)$, where $M$ is the $n^{k}+k$ time-bounded nondeterministic Turing machine itself and $k$ is the degree of polynomial $|x|^{k}+k$ such that $M(x)$ will halt within $|x|^{k}+k$ steps for any input $x$ of length $|x|$. We call such a positive integer $k$ the order of $(M,k)$.
By Lemma 2.7 (respectively, Lemma 2.8) stated above, we only need to discuss the single-tape nondeterministic Turing machines (respectively, single-tape ${\rm co}\mathcal{NP}$ machines). Thus, in the following context, by a nondeterministic Turing machine, we mean a single-tape nondeterministic Turing machine. Similarly, by a ${\rm co}\mathcal{NP}$ machine we mean a single-tape ${\rm co}\mathcal{NP}$ machine.
To obtain our main result, we must enumerate all ${\rm co}\mathcal{NP}$ machines and suppose that the set of all ${\rm co}\mathcal{NP}$ machines is enumerable (which needs to be proved) so that we can refer to the $j$-th ${\rm co}\mathcal{NP}$ machine in the enumeration. But before doing that, we should enumerate all polynomial-time nondeterministic Turing machines first, or, equivalently, prove that the set of all polynomial-time nondeterministic Turing machines is enumerable.
In what follows, we first use the method presented in [AHU74], p. 407, to encode a single-tape nondeterministic Turing machine into an integer.
Without loss of generality, we can make the following assumptions about the representation of a single-tape nondeterministic Turing machine with input alphabet $\{0,1\}$ because that will be all we need:
  1. (1)
    The states are named
    $q_{1},q_{2},\cdots,q_{s}$
    for some $s$, with $q_{1}$ the initial state and $q_{s}$ the accepting state.
  2. (2)
    The input alphabet is $\{0,1\}$.
  3. (3)
    The tape alphabet is
    $\{X_{1},X_{2},\cdots,X_{t}\}$
    for some $t$, where $X_{1}=\mathbbm{b}$, $X_{2}=0$, and $X_{3}=1$.
  4. (4)
    The next-move function $\delta$ is a list of quintuples of the form,
    $\{(q_{i},X_{j},q_{k},X_{l},D_{m}),\cdots,(q_{i},X_{j},q_{f},X_{p},D_{z})\}$
    meaning that
    $\delta(q_{i},X_{j})=\{(q_{k},X_{l},D_{m}),\cdots,(q_{f},X_{p},D_{z})\},$
    and $D_{m}$ is the direction, $L$, $R$, or $S$, if $m=1,2$, or $3$, respectively. We assume this quintuple is encoded by the string
    $110^{i}10^{j}10^{k}10^{l}10^{m}1\cdots 10^{i}10^{j}10^{f}10^{p}10^{z}$ 1 .
  5. (5)
    The nondeterministic Turing machine itself is encoded by concatenating in any order the codes for each of the quintuples in its next-move function. Additional $1$’s may be prefixed to the string if desired. The result will be some string of 00’s and $1$’s, beginning with $1$, which we can interpret as an integer.
Next, we encode the order of $(M,k)$ to be
$110^{k}$ 1
so that the tuple $(M,k)$ should be the concatenation of the binary string representing $M$ itself followed by the order $110^{k}$ 1. Now the tuple $(M,k)$ is encoded as a binary string, which can be explained as an integer.
By this encoding, any integer that cannot be decoded is assumed to represent the trivial Turing machine with an empty next-move function. Every single-tape polynomial-time nondeterministic Turing machine will appear infinitely often in the enumeration since, given a polynomial-time nondeterministic Turing machine, we may prefix $1$’s at will to find larger and larger integers representing the same set of $(M,k)$. We denote such a polynomial-time nondeterministic Turing machine by $\widehat{M}_{j}$, where $j$ is the integer representing the tuple $(M,k)$, i.e., $j$ is the integer value of the binary string representing the tuple $(M,k)$.
For convenience, we denote by $\{(M,k)\}$ the set of all polynomial-time nondeterministic Turing machines. Then, from the above arguments, we in fact have defined a $(1,1)$ correspondence between the set $\{(M,k)\}$ of all polynomial-time nondeterministic Turing machines and $\mathbb{N}_{1}$ if any integer that cannot be decoded is deemed to represent the trivial polynomial-time nondeterministic Turing machine. In other words, the set $\{(M,k)\}$ of all polynomial-time nondeterministic Turing machines is therefore enumerable. We further denote such an enumeration by $e$.
Therefore, we have reached the following important result:
Theorem 3.1.
There exists a one-one correspondence between $\mathbb{N}_{1}$ and the set $\{(M,k)\}$ of all polynomial-time nondeterministic Turing machines. In other words, the set $\{(M,k)\}$ of all polynomial-time nondeterministic Turing machines is enumerable (i.e., all polynomial-time nondeterministic Turing machines are in the list of $e$), and all languages in $\mathcal{NP}$ are enumerable (with languages appearing multiple times).
Remark 3.1.
There is another way to enumerate all polynomial-time nondeterministic Turing machines without encoding the order of polynomial into their representation. To do so, we need the Cantor pairing function (see Fig. 2 below from [A8]):
$\pi:\mathbb{N}\times\mathbb{N}\rightarrow\mathbb{N}$
defined by
$\pi(k_{1},k_{2}):=\frac{1}{2}(k_{1}+k_{2})(k_{1}+k_{2}+1)+k_{2}$
where $k_{1},k_{2}\in\mathbb{N}$. Since the Cantor pairing function is invertible (see [A8]), it is a bijection between $\mathbb{N}\times\mathbb{N}$ and $\mathbb{N}$. Note that we can further define a bijection $\tau$ between $\mathbb{N}$ and $\mathbb{N}_{1}$ such as $\tau:k\mapsto k+1$ where $k\in\mathbb{N}$ and $k+1\in\mathbb{N}_{1}$, which shows that there exists a $(1,1)$ correspondence between $\mathbb{N}$ and $\mathbb{N}_{1}$. As we have shown that any polynomial-time nondeterministic Turing machine is an integer, we can place any polynomial-time nondeterministic Turing machine and its order of polynomial into a tuple $(M,k)$ and use the Cantor pairing function to map the tuple $(M,k)$ to an integer in $\mathbb{N}_{1}$. The reader can easily check that such a method gives an enumeration of the set of polynomial-time nondeterministic Turing machines.
Figure 2. Cantor pairing function
In fact, by Definition 2.4, the above enumeration $e$ also gives an enumeration of all ${\rm co}\mathcal{NP}$ machines. To see this, we just force the non-trivial polynomial-time nondeterministic Turing machine in the enumeration $e$ to work in the condition of “there exists no accepting path.” Then, it is easy to see that each polynomial-time nondeterministic Turing machine corresponds to a ${\rm co}\mathcal{NP}$ machine, and vice versa.
Thus, from the above arguments, we can present the following important theorem.
Theorem 3.2.
The set of all ${\rm co}\mathcal{NP}$ machines is enumerable (i.e., all ${\rm co}\mathcal{NP}$ machines are enumerable). In other words, all languages in ${\rm co}\mathcal{NP}$ are enumerable (with languages appearing multiple times), i.e., all ${\rm co}\mathcal{NP}$ machines are in the list of $e$.
In the following context, we often use the notation $\langle M\rangle$ to denote the shortest binary string that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$. We should further remark that the shortest binary string $\langle M\rangle$ that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$ is by concatenating in any order the codes for each of the quintuples in its next-move function followed by the order $110^{k}$ 1 without additional $1$ prefixed to it, from which it is clear that $\langle M\rangle$ may not be unique (since the order of codes of the quintuples may be different), but the lengths of all different $\langle M\rangle$ are the same.
4. Simulation of All Polynomial-time Nondeterministic Turing Machines
In computational complexity, the most basic approach to show hierarchy theorems uses simulation and diagonalization, because the simulation and diagonalization technique [Tur37, HS65] is a standard method to prove lower bounds on uniform computing models (i.e., the Turing machine model); see for example [AHU74]. These techniques work well in deterministic time and space measures; see e.g. [For00, FS07]. However, they do not work for nondeterministic time, which is not known to be closed under complement at present — one of the main issues discussed in this paper; hence it is unclear how to define a nondeterministic machine that “does the opposite”; see e.g. [FS07].
For nondeterministic time, we can still do a simulation but can no longer negate the answer directly. In this case, we apply the lazy diagonalization to show the nondeterministic time hierarchy theorem; see e.g. [AB09]. It is worth noting that Fortnow [For11] developed a much more elegant and simple style of diagonalization to show nondeterministic time hierarchy; see also [FS10]. Generally, lazy diagonalization [Žák83, AB09, FS07] is a clever application of the standard diagonalization technique [Tur37, HS65, AHU74]. The basic strategy of lazy diagonalization is that for any $n\leq j<m=n^{t(n)}$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $1^{j+1}$, accepting if $M_{i}$ accepts and rejecting if $M_{i}$ rejects. When $j=m$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $x$ deterministically, accepting if $M_{i}$ rejects and rejecting if $M_{i}$ accepts. Since $m=n^{t(n)}$, $M$ has enough time to perform the trivial exponential-time deterministic simulation of $M_{i}$ on input $1^{n}$; what we are doing is deferring the diagonalization step by linking $M$ and $M_{i}$ together on larger and larger inputs until $M$ has an input large enough that it can actually “do the opposite” deterministically; see e.g. [FS07] for details.
However, although all ${\rm co}\mathcal{NP}$ machines are proved to be enumerable and the enumeration is $e$ given in Section 3, it is hard to apply the standard diagonalization technique or the lazy-diagonalization technique from [Žák83, AB09, FS07] to create a language that is not in ${\rm co}\mathcal{NP}$. Instead, our main goal of this section is to apply simulation techniques rather than diagonalization techniques and the particular relations between $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by a ${\rm co}\mathcal{NP}$ machine $M$ to show that there is a language $L_{s}$ accepted by a universal nondeterministic Turing machine with the property that $L_{s}\notin{\rm co}\mathcal{NP}$.
Now we can show the following important theorem:
Theorem 4.1.
There exists a language $L_{s}$ accepted by a universal nondeterministic Turing machine $U$, but not by any ${\rm co}\mathcal{NP}$ machines. Namely, there is a language $L_{s}$ such that
$L_{s}\notin{\rm co}\mathcal{NP}.$
Proof 4.1.
Let $U$ be a four-tape nondeterministic Turing machine which operates as follows on an input string $x$ of length of $n$.
  1. (1)
    $U$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}$ machine $M_{l}$ for some $l$ (by Definition 2.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic Turing machine, i.e., $x$ represents the trivial Turing machine with an empty next-move function) then GOTO $5$, else determines $t$, the number of tape symbols used by $M_{l}$; $s$, its number of states; and $k$, its order. The third tape of $U$ can be used as “scratch” memory to calculate $t$.
  2. (2)
    Then $U$ lays off on its second tape $n$ blocks of
    $\lceil\log t\rceil$
    cells each, the blocks being separated by single cell holding a marker ##\##, i.e., there are
    $(1+\lceil\log t\rceil)n$
    cells in all. Each tape symbol occurring in a cell of $M_{l}$’s tape will be encoded as a binary number in the corresponding block of the second tape of $U$. Initially, $U$ places $M_{l}$’s input, in binary coded form, in the blocks of tape $2$, filling the unused blocks with the code for the blank.
  3. (3)
    On tape $3$, $U$ sets up a block of
    $\lceil(k+1)\log n\rceil$
    cells, initialized to all 00’s. Tape $3$ is used as a counter to count up to
    $n^{k+1}.$
  4. (4)
    By using nondeterminism, $U$ simulates $M_{l}$, using tape $1$, its input tape, to determine the moves of $M_{l}$ and using tape $2$ to simulate the tape of $M_{l}$. The moves of $U$ are counted in binary in the block of tape $3$, and tape $4$ is used to hold the states of $M_{l}$. If the counter on tape $3$ overflows, $U$ halts without accepting. The specified simulation is the following:
    “On input $x=1^{n-m}\langle M_{l}\rangle$ where $m=|\langle M_{l}\rangle|$, then $U$ on input $1^{n-m}\langle M_{l}\rangle$ simulates $M_{l}$ on input $1^{n-m}\langle M_{l}\rangle$ using nondeterminism in
    $n^{k+1}$
    time and output its answer, i.e., accepting if $M_{l}$ accepts and rejecting if $M_{l}$ rejects.”
    It should be pointed out that, although we say that input $x$ is decoded into a ${\rm co}\mathcal{NP}$ machine $M_{l}$, $U$ in fact simulates the $n^{k}+k$ time-bounded nondeterministic Turing machine $M_{l}$. That is to say, $U$ accepts the input $x=1^{n-m}\langle M_{l}\rangle$ if there exists at least one accepting path of $M_{l}$ on input $x$, and $U$ rejects the input $x=1^{n-m}\langle M_{l}\rangle$ if there exists no accepting path of $M_{l}$ on input $x$.
  5. (5)
    Since $x$ is encoding of the trivial Turing machine with an empty next-move function. Then $U$ sets up a block of
    $\lceil 2\times\log n\rceil$
    cells on tape $3$, initialized to all 00’s. Tape $3$ is used as a counter to count up to
    $n^{2}.$
    By using its nondeterministic choices, $U$ moves as per the path given by $x$. The moves of $U$ are counted in binary in the block of tape $3$. If the counter on tape $3$ overflows, then $U$ halts. $U$ accepts $x$ if and only if there is a computation path from the start state of $U$ leading to the accept state and the total number of moves can not exceed
    $n^{2}$
    steps, so is within $O(n)$. Note that the number of $2$ in
    $\lceil 2\times\log n\rceil$
    is fixed, i.e., it is a default setting (or it is by default).
The nondeterministic Turing machine $U$ constructed above is of time complexity, say $S$, which is currently unknown. By Lemma 2.7, $U$ is equivalent to a single-tape nondeterministic $O(S^{2})$ time-bounded Turing machine, and it, of course, accepts some language $L_{s}$.
We claim that $L_{s}\notin{\rm co}\mathcal{NP}$. Indeed, suppose for the sake of contradiction that $L_{s}$ is decided by some $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}$ machine $M_{i}$. Then by Lemma 2.8, we may assume that $M_{i}$ is a single-tape $T(n)=n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}$ machine. By Definition 2.4, we can regard it as an $n^{k}+k$ time-bounded nondeterministic Turing machine, and we can further let $U$ on input $1^{n-m}\langle M_{i}\rangle$ simulate the $n^{k}+k$ time-bounded nondeterministic Turing machine $M_{i}$ on input $1^{n-m}\langle M_{i}\rangle$. Let $M_{i}$ have $s$ states and $t$ tape symbols, and the shortest length of $M_{i}$ is $m$, i.e.,
$m=\min\{|\langle M_{i}\rangle|\,:\,\text{different binary strings $\langle M_{i}\rangle$ represent the same $M_{i}$}\}.$
Since $M_{i}$ is represented by infinitely many strings and appears infinitely often in the enumeration $e$, $U$ was to set the counter of tape $3$ to count up to $n^{k+1}$, by Lemma 2.10, to simulate
$T(n)$
steps of $M_{i}$ (Note that the input of $M_{i}$ is $1^{n-m}\langle M_{i}\rangle$ where $m=|\langle M_{i}\rangle|$, so the length of $1^{n-m}\langle M_{i}\rangle$ is $n$), $U$ requires to perform
$T(n)\log T(n)$
steps, and since
$\lim_{n\rightarrow\infty}$ $\frac{T(n)\log T(n)}{n^{k+1}}$
$=$ $\lim_{n\rightarrow\infty}\frac{(n^{k}+k)\log(n^{k}+k)}{n^{k+1}}$
$=$ $\lim_{n\rightarrow\infty}\left(\frac{n^{k}\log(n^{k}+k)}{n^{k+1}}+\frac{k\log(n^{k}+k)}{n^{k+1}}\right)$
$=$ 00 0
$<$ $1$
So, there exists an $N_{0}>0$ such that for any $N>N_{0}$,
$T(N)\log T(N)<N^{k+1}$
which implies that for a sufficiently long $w$, say $|w|\geq N_{0}$, and $M_{w}$ denoted by such $w$ is $M_{i}$, we have
$T(|w|)\log T(|w|)<|w|^{k+1}.$
Thus, on input $w$, $U$ has sufficient time to simulate $M_{w}$. This yields that
(4.1)4.1( 4.1 ) $1^{|w|-m}\langle M_{i}\rangle\in L_{s}\quad\text{if and only if}\quad 1^{|w|-m}\langle M_{i}\rangle\in L(M_{w})=L(M_{i}).$
But by the relation between the languages $L(M_{i})$ accepted by polynomial-time nondeterministic Turing machine $M_{i}$ and $\overline{L}(M_{i})$ accepted by ${\rm co}\mathcal{NP}$ machine $M_{i}$, we have
(4.2)4.2( 4.2 ) $1^{|w|-m}\langle M_{i}\rangle\in L(M_{i})\quad\text{if and only if}\quad 1^{|w|-m}\langle M_{i}\rangle\notin\overline{L}(M_{i}).$
From (4.1) and (4.2) we conclude that if $U$ accepts the input $1^{|w|-m}\langle M_{i}\rangle$, i.e.,
$1^{|w|-m}\langle M_{i}\rangle\in L_{s}$
then
$1^{|w|-m}\langle M_{i}\rangle\notin\overline{L}(M_{i}),$
i.e.,
$\chi_{\overline{L}(M_{i})}(1^{|w|-m}\langle M_{i}\rangle)=0,$
and if $U$ rejects the input $1^{|w|-m}\langle M_{i}\rangle$, i.e.,
$1^{|w|-m}\langle M_{i}\rangle\notin L_{s}$
then
$1^{|w|-m}\langle M_{i}\rangle\in\overline{L}(M_{i}),$
i.e.,
$\chi_{\overline{L}(M_{i})}(1^{|w|-m}\langle M_{i}\rangle)=1,$
a contradiction with our assumption that $L_{s}$ is decided by the $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}$ machine $M_{i}$.
The above arguments further yield that there exists no ${\rm co}\mathcal{NP}$ machine $M_{i}$ in the enumeration $e$ accepting the language $L_{s}$. Since all ${\rm co}\mathcal{NP}$ machines are in the list of $e$, thus,
$L_{s}\notin{\rm co}\mathcal{NP}.$
This finishes the proof.
Remark 4.2.
One may wish to apply simulation techniques to the case of $\mathcal{P}$ versus ${\rm co}\mathcal{P}$ and doubt that simulation techniques are unable to produce a language not in ${\rm co}\mathcal{P}$. In this scenario, we remark that the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}$ via a universal deterministic Turing machine $U^{\prime}$, because a polynomial-time deterministic Turing machine ($\mathcal{P}$ machine) that works under the “there exists no accepting path” condition is still a polynomial-time deterministic Turing machine. Namely, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic Turing machines are symmetrical. Specifically, for an arbitrary deterministic Turing machine $M$, if $M$ accepts the input $w$, then there is only one accepting path, and if $M$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic Turing machines. In short, if we apply the same technique, i.e., the simulation technique, to create a language not in ${\rm co}\mathcal{P}$, then the universal deterministic Turing machine $U^{\prime}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{\prime})\in\mathcal{P}$ (the proof is similar to the case in Section 5 below). But by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic Turing machines, we clearly have that $L(U^{\prime})\in{\rm co}\mathcal{P}$. Thus, we are unable to claim that $L(U^{\prime})\notin{\rm co}\mathcal{P}$.
5. Proving That $L_{s}\in\mathcal{NP}$
The previous section aims to establish a lower bounds result, i.e., to show that there is a language $L_{s}$ not in ${\rm co}\mathcal{NP}$. We now turn our attention to the upper bounds, i.e., to show that the language $L_{s}$ accepted by the universal nondeterministic Turing machine $U$ is in fact in $\mathcal{NP}$.
As a matter of fact, the technique to establish the desired upper bounds theorem is essentially the same as the one developed in the author’s recent work [Lin21], because the discussed issue to establish desired upper bounds is essentially the same as [Lin21]. But the work to establish such upper bounds for the first time is very difficult, and to see why, see e.g. [Lin21].
Now, for our goals, we are going to show firstly that the universal nondeterministic Turing machine $U$ runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$:
Theorem 5.1.
The universal nondeterministic Turing machine $U$ constructed in the proof of Theorem 4.1 runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$.
Proof 5.1.
The simplest way to show the theorem is to prove that for any input $w$ to $U$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $U$ runs at most
$|w|^{i_{w}+1}$
steps, which can be done as follows.
On the one hand, if the input $x$ encodes a $n^{k}+k$ time-bounded nondeterministic Turing machine, then $U$ turns itself off mandatorily within
$|x|^{k+1}$
steps by the construction, so the corresponding integer $i_{x}$ is $k$ in this case (i.e., $i_{x}=k$). This holds true for all polynomial-time nondeterministic Turing machines as input with $k$ to be the order of that corresponding polynomial-time nondeterministic Turing machine.
But on the other hand, if the input $x$ does not encode some polynomial-time nondeterministic Turing machine, then the running time of $U$ is within time
$O(|x|)$
because $U$ must turn itself off when the counter exceeds
$|x|^{2}$
steps by the construction, so the corresponding integer $i_{x}$ is $1$ in this case (i.e., $i_{x}=1$).
In both cases we have show that for any input $w$ to $U$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $U$ runs at most $|w|^{i_{w}+1}$ steps. Thus, $U$ is a nondeterministic
$O(n^{k})$
time-bounded Turing machine for any $k\in\mathbb{N}_{1}$. By Lemma 2.7, there is a single-tape nondeterministic Turing machine $U^{\prime}$ equivalent to $U$ and $U^{\prime}$ runs within time
$O(S(n)^{2})=O(n^{2k})$
for any $k\in\mathbb{N}_{1}$.
The following theorem is our main result of this section describing the upper bounds result:
Theorem 5.2.
The language $L_{s}$ is in $\mathcal{NP}$ where $L_{s}$ is accepted by the universal nondeterministic Turing machine $U$ constructed in the proof of Theorem 4.1.
Proof 5.2.
Let us first define the family of languages
$\left\{L_{s}^{i}\right\}_{i\in\mathbb{N}_{1}}$
as follows:
$L_{s}^{i}\overset{{\rm def}}{=}$ language accepted by $U$ running within time $O(n^{i})$ for fixed $i\in\mathbb{N}_{1}$.
That is, $U$ turns itself off mandatorily when its moves made by itself
during the computation exceed $n^{i+1}$ steps,
which technically can be done by adding a new tape to $U$ as a counter to count up to
$n^{i+1}$
for a fixed $i\in\mathbb{N}_{1}$, meaning that $U$ turns itself off when the counter of tape $3$ exceeds
$n^{k+1}$
or the counter of the newly added tape exceeds
$n^{i+1}.$
Obviously, for each $i\in\mathbb{N}_{1}$, $L_{s}^{i}$ is a truncation of $L_{s}$.
Then by the construction of $U$, namely, for an arbitrary input $w$ to $U$, there is a corresponding integer $i_{w}\in\mathbb{N}_{1}$ such that $U$ runs at most
$|w|^{i_{w}+1}$
steps (in other words, $U$ runs at most $n^{i+1}$ steps for any $i\in\mathbb{N}_{1}$ where $n$ is the length of input, see Theorem 5.1 above), we have
(5.1)5.1( 5.1 ) $L_{s}=\bigcup_{i\in\mathbb{N}_{1}}L_{s}^{i}.$
Furthermore,
$L_{s}^{i}\subseteq L_{s}^{i+1},\quad\text{for each fixed $i\in\mathbb{N}_{1}$}$
since for any word $w\in L_{s}^{i}$ accepted by $U$ within $O(n^{i})$ steps, it surely can be accepted by $U$ within $O(n^{i+1})$ steps, i.e.,
$w\in L_{s}^{i+1}.$
This gives that for any fixed $i\in\mathbb{N}_{1}$,
(5.2)5.2( 5.2 ) $L_{s}^{1}\subseteq L_{s}^{2}\subseteq\cdots\subseteq L_{s}^{i}\subseteq L_{s}^{i+1}\subseteq\cdots$
Note further that for any fixed $i\in\mathbb{N}_{1}$, $L_{s}^{i}$ is the language accepted by the nondeterministic Turing machine $U$ within time $O(n^{i})$, i.e., at most $n^{i+1}$ steps, we thus obtain
(5.3)5.3( 5.3 ) $L_{s}^{i}\in{\rm NTIME}[n^{i}]\subseteq\mathcal{NP},\quad\text{for any fixed $i\in\mathbb{N}_{1}$}.$
Now, (5.1), together with (5.2) and (5.3) easily yields
$L_{s}\in\mathcal{NP},$
as desired.
Proof of Theorem 1.1. It is clear that Theorem 1.1 follows immediately from Theorem 4.1 and Theorem 5.2. This completes the proof.
Remark 5.3.
By Lemma 2.7 and Theorem 5.2, the single-tape nondeterministic Turing machine $U_{1}$ that is equivalent to the nondeterministic Turing machine $U$ (i.e., $L(U_{1})=L(U)$) is also in the enumeration $e$. Or when working in the “there exists no accepting path” condition, the single-tape ${\rm co}\mathcal{NP}$ machine $U_{1}$ is also in the enumeration $e$. Specifically, we know from the proof of Theorem 5.2 that, for any $i\in\mathbb{N}_{1}$, we can make the multi-tape universal nondeterministic Turing machine $U$ run at most $|w|^{i+1}$ steps for all input $w\in\{0,1\}^{*}$ (i.e., $U$ runs within time $O(n^{i+1})$); by Lemma 2.7, such a machine is equivalent to a single-tape nondeterministic $O(n^{2(i+1)})$ time-bounded Turing machine $U_{i}$, i.e., $L_{s}^{i}=L(U_{i})$, and the single-tape nondeterministic $n^{2(i+1)}+2(i+1)$ time-bounded Turing machine $(U_{i},2(i+1))$ is in the enumeration $e$. But we have that $L_{s}\neq\overline{L}(U)$ (note that $L_{s}=L(U)$). Indeed, since for any input $w\in\{0,1\}^{*}$, $w\in L_{s}$ if and only if $w\not\in\overline{L}(U)$.
By the proof of Theorem 5.2, we close this section by presenting the following corollary:
Corollary 5.3.
For each fixed $k\in\mathbb{N}_{1}$, it holds that
$L_{d}^{k}=\bigcup_{i\leq k}L_{d}^{i}\in\mathcal{NP}$
where $i\in\mathbb{N}_{1}$.
Proof 5.4.
It follows clearly from the relations (5.2) and (5.3).
6. Breaking the Relativization Barrier
The computation model we use in this section is the query machines, or the oracle Turing machines, which is an extension of the multi-tape Turing machine, i.e., Turing machines that are given access to a black box or “oracle” that can magically solve the decision problem for some language
$X\subseteq\{0,1\}^{*}.$
The machine has a special oracle tape on which it can write a string
$w\in\{0,1\}^{*}$
and in one step gets an answer to a query of the form
[…]
which can be repeated arbitrarily often with different queries. If $X$ is a difficult language (say, one that cannot be decided in polynomial time, or is even undecidable), then this oracle gives the Turing machine additional power. We first quote its formal definition as follows:
Definition 6.1 (cf. the notion of deterministic oracle Turing machines in [AB09]).
A nondeterministic oracle Turing machine is a nondeterministic Turing machine $M$ that has a special read-write tape we call $M$’s oracle tape and three special states $q_{query}$, $q_{yes}$, $q_{no}$. To execute $M$, we specify in addition to the input a language $X\subseteq\{0,1\}^{*}$ that is used as the oracle for $M$. Whenever during the execution $M$ enters the state $q_{query}$, the machine moves into the state $q_{yes}$ if $w\in X$ and $q_{no}$ if $w\not\in X$, where $w$ denotes the contents of the special oracle tape. Note that, regardless of the choice of $X$, a membership query to $X$ counts only as single computation step. If $M$ is an oracle machine, $X\subseteq\{0,1\}^{*}$ a language, and $x\in\{0,1\}^{*}$, then we denote the output of $M$ on input $x$ and with oracle $X$ by $M^{X}(x)$. An input $x$ is said to be accepted by a nondeterministic oracle Turing machine $M^{X}$ if there is a computation path of $M^{X}$ on input $x$ leading to the accepting state of $M^{X}$.
Next, we define the polynomial-time nondeterministic oracle Turing machines.
Definition 6.2.
Formally, a polynomial-time nondeterministic oracle Turing machine $M^{O}$ with oracle $O$ is a nondeterministic oracle Turing machine with oracle $O$ such that there exists $k\in\mathbb{N}_{1}$, for all input $w$ of length $|w|$ where $|w|\in\mathbb{N}$, $M^{O}(w)$ will halt within $t(|w|)=|w|^{k}+k$ steps.
Similarly, by default, a word $w$ is accepted by a nondeterministic $t(n)$ time-bounded oracle Turing machine $M^{X}$ with oracle $X$ if there exists at least one computation path $\gamma$ putting the machine into an accepting state. This is the “there exists at least one accepting path” condition for nondeterministic oracle Turing machines, based on which the complexity class ${\rm NTIME}^{X}[t(n)]$ is defined:
Definition 6.3 (“there exists at least one accepting path” condition).
The set of languages decided by nondeterministic oracle Turing machines with oracle $X$ within time $t(n)$ by the condition of “there exists at least one accepting path” is denoted by ${\rm NTIME}^{X}[t(n)]$. Thus,
$\mathcal{NP}^{X}=\bigcup_{k\in\mathbb{N}_{1}}{\rm NTIME}^{X}[n^{k}].$
Also, there is another condition (i.e., there exists no computation path $\gamma$ putting the machine into an accepting state) for nondeterministic oracle Turing machines, which is the dual of the “there exists at least one accepting path” condition, based on which the notion of oracle ${\rm co}\mathcal{NP}$ machine is defined. Namely, the following
Definition 6.4.
A ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}$ with oracle $A$ is exactly a polynomial-time nondeterministic oracle Turing machine $M^{A}$ with oracle $A$ working under the condition of “there exists no accepting path,” and we define the language accepted by ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}$ to be $\overline{L}(M^{A})$. Namely, for any input $w$, $w\in\overline{L}(M^{A})$ if and only if polynomial-time nondeterministic oracle Turing machine $M^{A}$ rejects $w$.
Similarly, for a ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}$ with oracle $A$, we define the characteristic function
$\chi_{\overline{L}(M^{A})}:\{0,1\}^{*}\rightarrow\{0,1\}$
as follows:
$\chi_{\overline{L}(M^{A})}(w)=\left\{\begin{array}[]{ll}1,&\hbox{$w\in\overline{L}(M^{A})$;}\\ 0,&\hbox{$w\notin\overline{L}(M^{A})$.}\end{array}\right.$
Let $M^{A}$ be a polynomial-time nondeterministic oracle Turing machine with oracle $A$. Since $M^{A}$ can be regarded as both a polynomial-time nondeterministic oracle Turing machine working under the condition of “there exists at least one accepting path” and a ${\rm co}\mathcal{NP}^{A}$ machine working under the condition of “there exists no accepting path”, we also strictly distinguish, in the following context, between these two situations in which the use of symbols. Namely, we use the notation
$M^{A}(w)=1\quad\text{(respectively, $M^{A}(w)=0$)}$
to mean that polynomial-time nondeterministic oracle Turing machine $M^{A}$ accepts (respectively, rejects) the input $w$; we use the notation
$\chi_{\overline{L}(M^{A})}(w)=1\quad\text{ (respectively, $\chi_{\overline{L}(M^{A})}(w)=0$)}$
to mean that ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}$ accepts (respectively, rejects) the input $w$.
It is clear that for all $w\in\{0,1\}^{*}$, $w\in\overline{L}(M^{A})$ if and only if
$M^{A}(w)=0,\quad\text{(i.e., polynomial-time nondeterministic oracle Turing machine $M^{A}$ rejects $w$)}$
and $w\notin\overline{L}(M^{A})$ if and only if
$M^{A}(w)=1,\quad\text{(i.e., polynomial-time nondeterministic oracle Turing machine $M^{A}$ accepts $w$).}$.
In a similar way to define the complexity class ${\rm coNTIME}[t(n)]$ in Section 2, we can define the complexity class ${\rm coNTIME}^{X}[t(n)]$, based on the “there exists no accepting path” condition, as follows.
Definition 6.5 (“there exists no accepting path” condition).
The set of languages decided by nondeterministic oracle Turing machines with oracle $X$ within time $t(n)$ by the condition of “there exists no accepting path” is denoted by ${\rm coNTIME}^{X}[t(n)]$, i.e., a $t(n)$ time-bounded nondeterministic oracle Turing machine with oracle $X$ accepts a string $w$ if and only if there exists no accepting path on input $w$. Thus,
${\rm co}\mathcal{NP}^{X}=\bigcup_{k\in\mathbb{N}_{1}}{\rm coNTIME}^{X}[n^{k}].$
In the following context, we will assume the notion of ${\rm co}\mathcal{P}^{X}$ machines, which can be defined in a similar way to that of Definition 6.4, and is omitted here.
In $1975$, Baker, Gill, and Solovay [BGS75] presented a proof of that:
There is an oracle A for which $\text{There is an oracle $A$ for which }\mathcal{P}^{A}=\mathcal{NP}^{A}.$
Baker, Gill, and Solovay [BGS75] suggested that their results imply that ordinary simulation and diagonalization techniques are not capable of proving $\mathcal{P}\neq\mathcal{NP}$. Likewise, if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, then Baker, Gill, and Solovay’s above result also indirectly suggests that ordinary simulation techniques are not capable of proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$.
To explore what is behind this kind of mystery, let us dive into two cases: one is $\mathcal{P}^{O}=\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O}$ for some oracle $O$, and another is $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ for some oracle $A$.
Note that, the result of Baker, Gill, and Solovay [BGS75] also implies that for the same oracle $O$,
$\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O},$
because if $\mathcal{P}^{O}=\mathcal{NP}^{O}$, then ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Further by
$\mathcal{P}^{O}={\rm co}\mathcal{P}^{O}.$
Thus, we have that
$\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O},$
which belongs to the first case mentioned above:
$\mathcal{P}^{O}=\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O}.$
It should be pointed out that, in this scenario, we are unable to apply simulation techniques to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$, for which the reasons will be given in Theorem 6.1 below.
Theorem 6.1.
If $\mathcal{P}^{O}=\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O}$, and assume that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, then simulation techniques are unable to create a language $L_{s}^{O}\notin{\rm co}\mathcal{NP}^{O}$, thus unable to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$.
Proof 6.6.
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$, thus we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, if the set of all ${\rm co}\mathcal{P}^{O}$ machines is not enumerable, then of course we can not apply simulation techniques. Thus, we assume that the set of ${\rm co}\mathcal{P}^{O}$ machines is enumerable, as the premise of the theorem states. Since we have assumed that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to the Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine that works under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.
Now, we assume another scenario, i.e., the case that there exists an oracle $A$ such that
$\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}.$
If the above case is true, let us explore what is behind this kind of mystery.
First of all, for convenience, let us denote the set of all ${\rm co}\mathcal{NP}^{X}$ machines by the notation $\mathcal{T}^{X}$.
Let us suppose that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$. In this case, to break the Relativization Barrier or to explore the mystery behind the implications of Baker, Gill, and Solovay’s result [BGS75], let us further make the following rational assumptions:
  1. (1)
    The oracle ${\rm co}\mathcal{NP}$ machines (by Definition 6.4, equivalently, the polynomial-time nondeterministic oracle Turing machines) can be encoded to a string over $\{0,1\}$;
  2. (2)
    There are universal nondeterministic oracle Turing machines that can simulate any other polynomial-time nondeterministic oracle Turing machine, which is similar to the behavior of the universal nondeterministic Turing machine $U$ constructed in Section 4;
  3. (3)
    The simulation can be done within in time
    $O(T(n)\log T(n)),$
    where $T(n)$ is the time complexity of the simulated polynomial-time nondeterministic oracle Turing machine.
Then, we will prove the following interesting theorems:
Theorem 6.2.
If $\mathcal{P}^{A}\neq\mathcal{NP}^{A}$ and suppose that the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines with oracle $A$ is enumerable and the enumeration is $e^{\prime}$, and we have the above rational assumptions. Then, there exists a language $L_{s}^{A}$ accepted by a universal nondeterministic oracle Turing machine $U^{A}$ with oracle $A$, but not by any ${\rm co}\mathcal{NP}^{A}$ machines with oracle $A$. Namely, there is a language $L_{s}^{A}$ such that
$L_{s}^{A}\notin{\rm co}\mathcal{NP}^{A}.$
Proof 6.7.
Since the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is enumerable, and apart from it, we further have the above rational Assumptions, i.e., the following: (1) The ${\rm co}\mathcal{NP}^{A}$ machine can be encoded to a string over $\{0,1\}$; (2) There exist universal nondeterministic oracle Turing machines with oracle $A$ that can simulate any other polynomial-time nondeterministic oracle Turing machine with oracle $A$, which is similar to the behavior of universal nondeterministic Turing machine $U$ constructed in Section 4; (3) The simulation can be done within time
$O(T(n)\log T(n)),$
where $T(n)$ is the time complexity of the simulated polynomial-time nondeterministic oracle Turing machine with oracle $A$. Then, the remainder of the proof is to show that there exists a universal nondeterministic oracle Turing machine $U^{A}$ with oracle $A$ accepting the language
$L_{s}^{A}\notin{\rm co}\mathcal{NP}^{A}$
in a similar way to that of Theorem 4.1. For convenience of the reader and for clarity, we proceed as follows: let $U^{A}$ be a five-tape universal nondeterministic oracle Turing machine that operates as follows on an input string $x$ of length of $n$:
  1. (1)
    $U^{A}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$ with oracle $A$ for some $j$ (by Definition 6.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic oracle Turing machine $M^{A}_{j}$ for some $j$, i.e., $x$ represents the trivial oracle Turing machine with an empty next-move function) then GOTO $6$, else determines $t$, the number of tape symbols used by $M^{A}_{j}$; $s$, its number of states; and $k$, its order. [^4] We suppose that the order $110^{k}$ 1 of $M^{A}_{j}$ for some $k\in\mathbb{N}_{1}$ is also encoded into the binary string representing ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$, which is similar to Section 3. The third tape of $U^{A}$ can be used as “scratch” memory to calculate $t$.
  2. (2)
    Then $U^{A}$ lays off on its second tape $n$ blocks of
    $\lceil\log t\rceil$
    cells each, the blocks being separated by single cell holding a marker ##\##, i.e., there are
    $(1+\lceil\log t\rceil)n$
    cells in all. Each tape symbol occurring in a cell of $M^{A}_{j}$’s tape will be encoded as a binary number in the corresponding block of the second tape of $U^{A}$. Initially, $U^{A}$ places $M^{A}_{j}$’s input, in binary coded form, in the blocks of tape $2$, filling the unused blocks with the code for the blank.
  3. (3)
    On tape $3$, $U^{A}$ sets up a block of
    $\lceil(k+1)\log n\rceil$
    cells, initialized to all 00’s. Tape $3$ is used as a counter to count up to
    $n^{k+1}.$
  4. (4)
    On tape $4$, $U^{A}$ reads and writes the contents of oracle tape of $M^{A}_{j}$. That is, tape $4$ is the oracle tape of $U^{A}$ which is used to simulate the oracle tape of $M^{A}_{j}$.
  5. (5)
    By using nondeterminism, $U^{A}$ simulates $M^{A}_{j}$, using tape $1$, its input tape, to determine the moves of $M^{A}_{j}$ and using tape $2$ to simulate the tape of $M^{A}_{j}$. The moves of $U^{A}$ are counted in binary in the block of tape $3$, and tape $5$ is used to hold the states of $M^{A}_{j}$. If the counter on tape $3$ overflows, $U^{A}$ halts without accepting. The specified simulation is the following:
    “On input $x=1^{n-m}\langle M^{A}_{j}\rangle$ where $m=|\langle M^{A}_{j}\rangle|$, then $U^{A}$ on input $1^{n-m}\langle M^{A}_{j}\rangle$ simulates $M^{A}_{j}$ on input $1^{n-m}\langle M^{A}_{j}\rangle$ using nondeterminism in
    $n^{k+1}$
    time and output its answer, i.e., accepting if $M^{A}_{j}$ accepts and rejecting if $M^{A}_{j}$ rejects.”
    Also note that, although we say that input $x$ is decoded into a ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$, $U^{A}$ in fact simulates the $n^{k}+k$ time-bounded nondeterministic oracle Turing machine $M^{A}_{j}$ with oracle $A$. Namely, $U^{A}$ accepts the input $x=1^{n-m}\langle M^{A}_{j}\rangle$ if there exists at least one accepting path of $M^{A}_{j}$ on input $x$, and $U^{A}$ rejects the input $x=1^{n-m}\langle M^{A}_{j}\rangle$ if there exists no accepting path of $M^{A}_{j}$ on input $x$.
  6. (6)
    Since $x$ is encoding of the trivial oracle Turing machine with an empty next-move function. Then $U^{A}$ sets up a block of
    $\lceil 2\times\log n\rceil$
    cells on tape $3$, initialized to all 00’s. Tape $3$ is used as a counter to count up to
    $n^{2}.$
    By using its nondeterministic choices, $U^{A}$ moves as per the path given by $x$. The moves of $U^{A}$ are counted in binary in the block of tape $3$. If the counter on tape $3$ overflows, then $U^{A}$ halts. $U^{A}$ accepts $x$ if and only if there is at least one computation path from the start state of $U^{A}$ leading to the accept state and the total number of moves can not exceed
    $n^{2}$
    steps, so is within
    $O(n)$
    steps. Note that the number of $2$ in
    $\lceil 2\times\log n\rceil$
    is fixed, i.e., it is a default setting.
The nondeterministic oracle Turing machine $U^{A}$ with oracle $A$ constructed above is of time complexity, say $S$, which is currently unknown. It, of course, accepts some language $L^{A}_{s}$.
We claim that $L^{A}_{s}\notin{\rm co}\mathcal{NP}^{A}$. Indeed, suppose for the sake of contradiction that $L^{A}_{s}$ is decided by some $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$. Then, we may assume that $M^{A}_{i}$ is a two-tape (with an input tape and an oracle tape) $T(n)=n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine. By Definition 6.4, we can regard it as an $n^{k}+k$ time-bounded nondeterministic oracle Turing machine with oracle $A$, and we can further let $U^{A}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$ simulate $M^{A}_{i}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$. Let $M^{A}_{i}$ have $s$ states and $t$ tape symbols and the shortest length of $M^{A}_{i}$ is $m$, i.e,
$m=\min\{|\langle M^{A}_{i}\rangle|\,:\,\text{different binary strings $\langle M^{A}_{i}\rangle$ represent the same $M^{A}_{i}$}\}.$
Since $M^{A}_{i}$ is represented by infinitely many strings and appears infinitely often in the enumeration $e^{\prime}$, and $U^{A}$ was to set the counter of tape $3$ to count up to $n^{k+1}$, by Assumption (3), to simulate
$T(n)$
steps of $M^{A}_{i}$, $U^{A}$ requires to perform
$T(n)\log T(n)$
steps, and since
$\lim_{n\rightarrow\infty}$ $\frac{T(n)\log T(n)}{n^{k+1}}$
$=$ $\lim_{n\rightarrow\infty}\frac{(n^{k}+k)\log(n^{k}+k)}{n^{k+1}}$
$=$ $\lim_{n\rightarrow\infty}\left(\frac{n^{k}\log(n^{k}+k)}{n^{k+1}}+\frac{k\log(n^{k}+k)}{n^{k+1}}\right)$
$=$ 00 0
$<$ $1$
So, there exists an $N^{\prime}>0$ such that for any $N>N^{\prime}$,
$T(N)\log T(N)<N^{k+1}$
which implies that for a sufficiently long $w$, say $|w|\geq N^{\prime}$, and $M^{A}_{w}$ denoted by such $w$ is $M^{A}_{i}$, we have
$T(|w|)\log T(|w|)<|w|^{k+1}.$
Thus, on input $w$, $U^{A}$ has sufficient time to simulate $M^{A}_{w}$. This yields that
(6.1)6.1( 6.1 ) $1^{|w|-m}\langle M^{A}_{i}\rangle\in L^{A}_{s}\quad\text{if and only if}\quad 1^{|w|-m}\langle M^{A}_{i}\rangle\in L(M^{A}_{w})=L(M^{A}_{i}).$
But by the relation between the languages $L(M^{A}_{i})$ accepted by polynomial-time nondeterministic orale Turing machine $M^{A}_{i}$ with oracle $A$ and $\overline{L}(M^{A}_{i})$ accepted by ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$, we have
(6.2)6.2( 6.2 ) $1^{|w|-m}\langle M^{A}_{i}\rangle\in L(M^{A}_{i})\quad\text{if and only if}\quad 1^{|w|-m}\langle M^{A}_{i}\rangle\notin\overline{L}(M^{A}_{i}).$
From (6.1) and (6.2) we conclude that if $U^{A}$ accepts the input $1^{|w|-m}\langle M^{A}_{i}\rangle$, i.e.,
$1^{|w|-m}\langle M^{A}_{i}\rangle\in L^{A}_{s}$
then
$1^{|w|-m}\langle M^{A}_{i}\rangle\notin\overline{L}(M^{A}_{i}),$
i.e.,
$\chi_{\overline{L}(M^{A}_{i})}(1^{|w|-m}\langle M^{A}_{i}\rangle)=0,$
and if $U^{A}$ rejects the input $1^{|w|-m}\langle M^{A}_{i}\rangle$, i.e.,
$1^{|w|-m}\langle M^{A}_{i}\rangle\notin L^{A}_{s}$
then
$1^{|w|-m}\langle M^{A}_{i}\rangle\in\overline{L}(M^{A}_{i}),$
i.e.,
$\chi_{\overline{L}(M^{A}_{i})}(1^{|w|-m}\langle M^{A}_{i}\rangle)=1,$
a contradiction with our assumption that $L^{A}_{s}$ is decided by the $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$.
The above arguments further yield that there exists no ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$ in the enumeration $e^{\prime}$ accepting the language $L^{A}_{s}$. Since all ${\rm co}\mathcal{NP}^{A}$ machines are in the list of $e^{\prime}$, thus,
$L^{A}_{s}\notin{\rm co}\mathcal{NP}^{A}.$
This finishes the proof.
Next, we are going to show that the universal nondeterministic oracle Turing machine $U^{A}$ with oracle $A$ runs within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$ as follows:
Theorem 6.3.
The universal nondeterministic oracle Turing machine $U^{A}$ with oracle $A$ constructed in the proof of Theorem 6.2 runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$.
Proof 6.8.
The proof of this theorem is similar to that of Theorem 5.1. Namely, to show the theorem is to prove that for any input $w$ to $U^{A}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $U^{A}$ runs at most
$|w|^{i_{w}+1}$
steps. Since the rest of the proof is similar to that of Theorem 5.1, the detail is thus omitted.
We are ready to prove the language $L^{A}_{s}$ is, in fact, in $\mathcal{NP}^{A}$:
Theorem 6.4.
The language $L_{s}^{A}$ accepted by the nondeterministic oracle Turing machine $U^{A}$ with oracle $A$ is in fact in $\mathcal{NP}^{A}$.
Proof 6.9.
The proof of this theorem is similar to that of Theorem 5.2. Namely, the first step is to define the family of languages
$\left\{L^{A,i}_{s}\right\}_{i\in\mathbb{N}_{1}}.$
Then, by proof of Theorem 6.3, we have that
(6.3)6.3( 6.3 ) $L_{s}^{A}=\bigcup_{i\in\mathbb{N}_{1}}L^{A,i}_{s}.$
Next, we must have to show that for all $i\in\mathbb{N}_{1}$,
(6.4)6.4( 6.4 ) $L^{A,i}_{s}\in\mathcal{NP}^{A},$
and
(6.5)6.5( 6.5 ) $L^{A,1}_{s}\subseteq L^{A,2}_{s}\subseteq\cdots\subseteq L^{A,i}_{s}\subseteq L^{A,i+1}_{s}\subseteq\cdots$
Finally, the conclusion follows from (6.3), (6.4), and (6.5). Since the rest of the proof is similar to that of Theorem 5.2, the more specific details are thus omitted.
Combining Theorem 6.2 and Theorem 6.4 actually yields the following result:
Theorem 6.5.
If $\mathcal{P}^{A}\neq\mathcal{NP}^{A}$. Assume that the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is enumerable. Further, under the above rational assumptions, we have
$\mathcal{NP}^{A}\neq{\rm co}\mathcal{NP}^{A}.$
But now we have the fact that $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, so our assumption in Theorem 6.5 is not true, i.e., the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable. Indeed, by $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, there exists a ${\rm co}\mathcal{NP}^{A}$ machine accepting the language $L^{A}_{s}$ that differs from any languages accepted by ${\rm co}\mathcal{NP}^{A}$ machines in the enumeration $e^{\prime}$. Thus, the ${\rm co}\mathcal{NP}^{A}$ machine accepting the language $L_{s}^{A}$ is not in the enumeration $e^{\prime}$, showing that the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is in fact not enumerable. So the truths behind this kind of mystery are the following theorem:
Theorem 6.6.
Under some rational assumptions (i.e., the above rational Assumptions (1), (2), and (3)), and if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, then the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable. Thereby, ordinary simulation techniques will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
Now, we are naturally at a right point to present the proof of Theorem 1.2 as follows: Proof of Theorem 1.2. Note that the statement of Theorem 1.2 is precisely the same as Theorem 6.6, thereby we have done the proof.
Remark 6.10.
Thus, from Theorem 1.2, we know that Baker, Gill, and Solovay’s result [BGS75] implies that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ is not necessarily a necessary and sufficient condition for $\mathcal{NP}^{O}\neq{\rm co}\mathcal{NP}^{O}$ for all oracle $O$. Likewise, Yao’s conclusion [Yao85] on oracle separation of polynomial hierarchy also implies that $\mathcal{NP}={\rm co}\mathcal{NP}$ is not necessarily a necessary and sufficient condition for $\mathcal{NP}^{O}={\rm co}\mathcal{NP}^{O}$ for all oracle $O$.
7. Rich Structure of ${\rm co}\mathcal{NP}$
In complexity theory, or computational complexity, problems that are in the complexity class $\mathcal{NP}$ but are neither in the class $\mathcal{P}$ nor $\mathcal{NP}$-complete are called $\mathcal{NP}$-intermediate, and the class of such problems is called $\mathcal{NPI}$. The well-known Ladner’s theorem, proved in 1975 by Ladner [Lad75], is a result asserting that if the complexity classes $\mathcal{P}$ and $\mathcal{NP}$ are different, then the class $\mathcal{NPI}$ is not empty. Namely, the complexity class $\mathcal{NP}$ contains problems that are neither in $\mathcal{P}$ nor $\mathcal{NP}$-complete. We call such a nice theorem the rich structure of the complexity class $\mathcal{NP}$.
Our main goal in this section is to establish a result saying that similar to the complexity class $\mathcal{NP}$ having a rich structure, the complexity class ${\rm co}\mathcal{NP}$ also has intermediate languages that are not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete. To do so in a simple way, we need to quote the following useful result whose proof can be found in [Lad75]:
Lemma 7.1 ([Lad75]).
(Suppose that $\mathcal{P}\neq\mathcal{NP}$). There is a language $L_{\rm inter}\in\mathcal{NP}$ such that $L_{\rm inter}$ is not in $\mathcal{P}$ and $L_{\rm inter}$ is not $\mathcal{NP}$-complete.
Next, we need to show that the language of all tautologies is ${\rm co}\mathcal{NP}$-complete:
Lemma 7.2.
Let ${\rm TAUT}=\{\varphi\,:\,\varphi\text{ is a tautology}\}$. Then the language ${\rm TAUT}$ is ${\rm co}\mathcal{NP}$-complete.
Proof 7.3.
It is sufficient to show that for every language $L\in{\rm co}\mathcal{NP}$, $L\leq_{p}{\rm TAUT}$ where $\leq_{p}$ is polynomial-time many-one reduction [Kar72]. We can modify the Cook-Levin reduction [Coo71, Lev73] from $\overline{L}=\{0,1\}^{*}-L$ to ${\rm SAT}$ (recall that all nondeterministic Turing machines are with input alphabet $\{0,1\}$ because that will be all we need, see Section 3). So for any input $w\in\{0,1\}^{*}$ the Cook-Levin reduction produces a formula $\varphi_{w}$ that is satisfiable if and only if $w\in\overline{L}$. In other words, the formula $\neg\varphi_{x}$ is in ${\rm TAUT}$ if and only if $x\in L$, which completes the proof.
Now, we are naturally at a right point to give the proof of Theorem 1.6 as follows:
Proof of Theorem 1.6. Let $L=\{0,1\}^{*}-L_{\rm inter}$, where the language $L_{\rm inter}$ is the same as in Lemma 7.1. Such a language indeed exists by Corollary 1.3, which states that $\mathcal{P}\neq\mathcal{NP}$.
Then $L\in{\rm co}\mathcal{NP}$, and we will show that $L$ is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete. To see so, suppose that $L\in\mathcal{P}$; then since $\mathcal{P}={\rm co}\mathcal{P}$, we have $\overline{L}=L_{\rm inter}\in\mathcal{P}$, a contradiction to the fact that $L_{\rm inter}\notin\mathcal{P}$ by Lemma 7.1. Further assume that $L$ is ${\rm co}\mathcal{NP}$-complete, then ${\rm TAUT}\leq_{p}L$. By Lemma 7.2, $\overline{L}=L_{\rm inter}$ is $\mathcal{NP}$-complete; it is also a contradiction to the fact that $L_{\rm inter}$ is not $\mathcal{NP}$-complete. Thereby, to summarize from the above two cases, we have shown that $L$ is ${\rm co}\mathcal{NP}$-intermediate. This finishes the proof.
8. Frege Systems
The focal point of this section is Frege systems, which are very strong proof systems in a propositional setting [CR79], based on axiom schemes and rules such as modus ponens. While Frege systems operate with Boolean formulas as lines, the extended Frege system EF works with Boolean circuits; see e.g. [Jer05]. Furthermore, showing lower bounds on Frege or even extended Frege systems constitutes a major open problem in proof complexity; see e.g. [BP01].
Informally, a proof system for a language $\mathcal{L}$ is a definition of what is considered to be a proof that $\Phi\in\mathcal{L}$; see e.g. [CR79]. The key features of a proof system are that it is sound, i.e., only formulas in $\mathcal{L}$ have proofs, and complete, i.e., all formulas in $\mathcal{L}$ have proofs, and that there is an algorithm with running time polynomial in $|\pi|$ to check whether $\pi$ is a proof that $\Phi\in\mathcal{L}$ [BH19].
Let $\mathfrak{F}$ be the set of functions $f:\Sigma_{1}^{*}\rightarrow\Sigma_{2}^{*}$, where $\Sigma_{1}$ and $\Sigma_{2}$ are any finite alphabets, such that $f$ can be computed by a deterministic Turing machine in time bounded by a polynomial in the length of the input. Then, we have the following definition:
Definition 8.1 ([CR79]).
If $L\subseteq\Sigma^{*}$, a proof system for $L$ is a function $f:\Sigma_{1}^{*}\rightarrow L$ for some alphabet $\Sigma_{1}$ and $f\in\mathfrak{F}$ such that $f$ is onto. We say that the proof system is polynomially bounded iff there is a polynomial $p(n)$ such that for all $y\in L$ there is a $x\in\Sigma_{1}^{*}$ such that $y=f(x)$ and $|x|\leq p(|y|)$, where $|z|$ denotes the length of a string $z$.
If $y=f(x)$, then $x$ is a proof of $y$, and $x$ is a short proof of $y$, if in addition, $|x|\leq p(|y|)$. Thus a proof system $f$ is polynomially bounded iff there is a bounding polynomial $p(n)$ with respect to which every $y\in L$ has a short proof.
In particular, Frege proof systems are proof systems for propositional logic. As a matter of fact, Frege systems are the usual “textbook” proof systems for propositional logic based on axioms and rules; see e.g. [CR79]. A Frege system composed of a finite set of axiom schemes and a finite number of rules is a possible axiom scheme. Furthermore, a Frege proof is a sequence of formulas where each formula is either a substitution instance of an axiom or can be inferred from previous formulas by a valid inference rule. At the same time, Frege systems are required to be sound and complete. However, the exact choice of the axiom schemes and rules does not matter, as any to Frege system are polynomial equivalent; see e.g.[CR79] or see Theorem 8.1 below. Thus, we can assume without loss of generality that modus ponens (see e.g. Remark 8.3 below) is the only rule of inference.
Definition 8.2 ([BG98]).
A Frege proof system $F$ is an inference system for propositional logic based on
  1. (1)
    a language of well-formed formulas obtained from a numerable set of propositional variables and any finite propositionally complete set of connectives;
  2. (2)
    a finite set of axiom schemes;
  3. (3)
    and the rule of Modus Ponens
    $\frac{A\quad A\rightarrow B}{B}.$
A proof $P$ of the formula $A$ in a Frege system is a sequence $A_{1},\cdots,A_{n}$ of formulas such that $A_{n}$ is $A$ and every $A_{i}$ is either an instance of an axiom scheme or it is obtained by the application of the Modus Ponens from premises $A_{j}$ and $A_{k}$ with $j,k<i$.
Remark 8.3.
The well-known inference rule of Modus Ponens is its only rule of inference:
$\frac{A\quad A\rightarrow B}{B}$
We give a more intuitive explanation about the inference rule of Modus Ponens: For example, in the proof of Theorem 5.2 in Section 5 where we prove that $L_{s}\in\mathcal{NP}$, $A$ is the following:
$A\overset{\rm def}{=}(5.1)\bigwedge(5.2)\bigwedge(5.3),$
$B$ represents the proposition $L_{s}\in\mathcal{NP}$. Namely,
$B\overset{\rm def}{=}L_{s}\in\mathcal{NP},$
and one of the main deductions of Theorem 5.2, besides showing that $A$ is true, is to show that $A$ implies $B$.
The axiom schemes of a Frege proof system is the following (see e.g. [Bus99]):
$(P\wedge Q)\rightarrow P$
$(P\wedge Q)\rightarrow Q$
$P\rightarrow(P\vee Q)$
$Q\rightarrow(P\vee Q)$
$(P\rightarrow Q)\rightarrow((P\rightarrow\neg Q)\rightarrow\neg P)$
($(\neg\neg P)\rightarrow P$
$P\rightarrow(Q\rightarrow P\wedge Q)$
$(P\rightarrow R)\rightarrow((Q\rightarrow R)\rightarrow(P\vee Q\rightarrow R))$
$P\rightarrow(Q\rightarrow P)$
$(P\rightarrow Q)\rightarrow(P\rightarrow(Q\rightarrow R))\rightarrow(P\rightarrow R).$
More generally, a Frege system is specified by any finite complete set of Boolean connectives and a finite set of axiom schemes and rule schemes, provided it is implicationally sound and implicationally complete.
Definition 8.4.
The length of a Frege proof is the number of symbols in the proof. The length $|\varphi|$ of a formula $\varphi$ is the number of symbols in $\varphi$.
Now, we introduce the notion of $p$-simulation between two proof systems.
Definition 8.5 ([CR79]).
If $f_{1}:\Sigma_{1}^{*}\rightarrow L$ and $f_{2}:\Sigma_{2}^{*}\rightarrow L$ are proof system for $L$, then $f_{2}$ $p$-simulates $f_{1}$ provided there is a function $g:\Sigma_{1}^{*}\rightarrow\Sigma_{2}^{*}$ such that $g\in\mathfrak{F}$, and $f_{2}(g(x))=f_{1}(x)$ for all $x$.
The notion of an abstract propositional proof system is given as follows:
Definition 8.6 ([Bus99]).
An abstract propositional proof system is a polynomial-time computable function $g$ such that $Range(g)=TAUT$; i.e., the range of $g$ is the set of all Boolean tautologies. An $g$-proof of a formula $\phi$ is a string $w$ such that $g(w)=\phi$.
8.1. Lower Bounds for Frege Systems
Let $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ be two arbitrary Frege systems, the following theorem indicates that these standard proof systems for the propositional calculus are about equally powerful:
Theorem 8.1 ([Rec76, CR79]).
Any two Frege systems $p$-simulate each other. Hence one Frege system is polynomially bounded if and only if all Frege systems are.
The following theorem gives a necessary and sufficient condition for the complexity class $\mathcal{NP}$ being closed under complementation. For completeness, the proof is also quoted from [CR79]:
Theorem 8.2 ([CR79]).
$\mathcal{NP}$ is closed under complementation if and only if $TAUT$ is in $\mathcal{NP}$.
Proof 8.7.
The “if” part. Suppose that the set of tautologies is in $\mathcal{NP}$. Then every set $L$ in $\mathcal{NP}$ is reducible to the complement of the tautologies [Coo71], i.e., there is a function $f$ that can be computed by a deterministic Turing machine in time bounded by a polynomial in the length of the input such that for all strings $x$, $x\in L$ iff $f(x)$ is not a tautology. Hence a nondeterministic procedure for accepting the complement of $L$ is: on input $x$, compute $f(x)$, and accept $x$ if $f(x)$ is a tautology, using the nondeterministic algorithm for tautologies assumed above. Hence the complement of $L$ is in $\mathcal{NP}$.
The “only if” part. Suppose that $\mathcal{NP}$ is closed under complementation. The the complement of the set of tautologies is in $\mathcal{NP}$, since to verify that a formula is not a tautology one can guess at a truth assignment and verify that it falsifies the formula.
Since nondeterministic Turing machines can simulate Frege proof (see e.g. [Bus99]) and by Theorem 8.1 we know that any two Frege systems $p$-simulate each others, we then can present the proof of Theorem 1.8 as follows:
Proof of Theorem 1.8. We show the theorem by contradiction. Suppose that there is a polynomial $p(n)$ such that for all $\psi\in TAUT$, there exists a proof $\oint$ of $\psi$ of length at most $p(|\psi|)$ (see Definition 8.4), then a nondeterministic Turing machine $M$ on input $\psi$ can guess $\oint$ of length at most $p(|\psi|)$ and check whether $\oint$ is correct in deterministic polynomial time, and we know that the set of tautologies is ${\rm co}\mathcal{NP}$-complete, which, together with Theorem 8.2, implies
$\mathcal{NP}={\rm co}\mathcal{NP},$
a contradiction to Theorem 1.1. This finishes the proof.
Remark 8.8.
Initially, propositional proof complexity has been primarily concerned with proving lower bounds (even conditional) for the length of proofs in propositional proof systems, which is extremely interesting and well justified in its own right, with the ultimate goal of settling whether $\mathcal{NP}={\rm co}\mathcal{NP}$ (see e.g. [CR79]). In fact, as mentioned in Section 1, such research directions are called Cook’s Program for separating $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (see e.g. [Bus12]): Prove superpolynomial lower bounds for proof lengths in stronger and stronger propositional proof systems until they are established for all abstract proof systems. However, our approach is “to do the opposite,” that is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems.
9. Open Problems
Does $\mathcal{P}=\mathcal{NP}\cap{\rm co}\mathcal{NP}$? Currently, we have no answer to this problem. We tried to think about it but failed. We remark that this is a fascinating open problem in complexity theory. It is easy to see that $\mathcal{P}\subseteq\mathcal{NP}\cap{\rm co}\mathcal{NP}$ and the real problem is whether $\mathcal{NP}\cap{\rm co}\mathcal{NP}\subseteq\mathcal{P}$. Note that an example of a problem that is known to belong to both $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (but not known to be in $\mathcal{P}$) is Integer factorization [RSA78]: given positive integers $m$ and $n$, determine if $m$ has a factor less than $n$ and greater than one.
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contains the problem Satisfiability? Note that by footnote $1$, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this can not exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. If one can show that Satisfiability is not in $\mathcal{BQP}$, then the relationship between $\mathcal{BQP}$ and $\mathcal{NP}$ were clear. Currently, we have no answer to this problem. We tried to think about it but also failed. We remark that this is also a fascinating open problem in complexity theory.
Appendix A Two Definitions of $\mathcal{NP}$ Are Equivalent
The official descriptions on the definitions of $\mathcal{NP}$ in [Coo00] is given by:
Definition A.1.
$\mathcal{NP}\overset{{\rm def}}{=}\{L\,:\,w\in L\Leftrightarrow\exists y(|y|\leq|w|^{k}\,\,\,\text{and }R(w,y))\}$
where the checking relation
$R(w,y)$
is polynomial-time, i.e., the language
$L_{R}\overset{{\rm def}}{=}\{w\#y\,:\,R(w,y)\}$
is in $\mathcal{P}$.
The complexity class $\mathcal{P}$, which is referred to by Definition A.1, is defined as follows:
Definition A.2.
$\begin{array}[]{ll}\mathcal{P}\overset{{\rm def}}{=}&\{L\,:\,L=L(M)\text{ for some Turing machine $M$ that runs}\\ &\qquad\text{ in polynomial time }\}.\end{array}$
In Definition A.2, Turing machine $M$ refers to a deterministic Turing machine, and the notion of “running in polynomial time” is that if there exists $k$ ($\in\mathbb{N}_{1}$) such that,
$T_{M}(n)\leq n^{k}+k,\quad\forall n\in\mathbb{N}_{1}$
where $T_{M}(n)$ is defined by
$T_{M}(n)=\max\{t_{M}(w)\,:\,w\in\Sigma^{n}\}$
and $t_{M}(w)$ denotes the number of steps (or moves) in the computation of $M$ on input $w$ of length $n$.
As you may note that, this definition (i.e., Definition A.1) looks a little different from that given in Section 2. But they are equivalent. Here, we will show the equivalence for these two definitions of $\mathcal{NP}$.
A.1. A Proof of Two Definitions of $\mathcal{NP}$ Are Equivalent.
The remainder of this Appendix is to show that Definition A.1 and the definition of $\mathcal{NP}$ given in Section 2 are equivalent.
To begin, we should first give another definition of $\mathcal{NP}$ which is similar to the official descriptions of the definition of $\mathcal{P}$ in [Coo00] (i.e., Definition A.2 above):
Definition A.3.
$\begin{array}[]{ll}\mathcal{NP}\overset{{\rm def}}{=}&\{L\,:\,L=L(M)\text{ forsome nondeterministic Turing machine $M$}\\ &\qquad\text{that runs in polynomial time}\}.\end{array}$
Now, let us finish the last step of the task that we mentioned at the beginning of this subsection.
Proof A.4.
We need to show that Definition A.3 is equivalent to Definition A.1 and to show the equivalence between Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2. By a simple argument, it is clear that Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2 are equivalent. We proceed as follows:
We show first the “if” part. Suppose that
$L\in\bigcup_{i\in\mathbb{N}_{1}}{\rm NTIME}[n^{i}].$
Then there is a $k\in\mathbb{N}_{1}$, such that
$L\in{\rm NTIME}[n^{k}]$
which means that for all $n\in\mathbb{N}_{1}$, there is a nondeterministic Turing machine $M$, for any $w\in\Sigma^{n}$
$T_{M}(|w|)\leq c_{0}n^{k}+c_{1}n^{k-1}+\cdots+c_{k-1}n+c_{k}\quad\text{where $c_{0}>0$}.$
and
$L=L(M).$
For any constants $c_{0}$, $c_{1}$, $\cdots$, $c_{k}$, there must exists a minimal $t\in\mathbb{N}_{1}$ such that for all $n\in\mathbb{N}$ and for any $w\in\Sigma^{n}$
$T_{M}(|w|)\leq n^{t}+t.$
So,
$L\in\{\mathcal{L}\,:\,\mathcal{L}=L(M)\text{ for some nondeterministic Turing machine $M$ that runs in polynomial time}\}.$
We show next the “only if” part. Suppose now that the language
$L\in\{\mathcal{L}\,:\,\mathcal{L}=L(M)\,\text{ for some nondeterministic Turing machine $M$ that runs in polynomial time}\}.$
Then there exists a $k\in\mathbb{N}_{1}$ such that for all $n\in\mathbb{N}_{1}$ and for all $w\in\Sigma^{n}$,
$T_{M}(|w|)\leq n^{k}+k$
which implies that
$L\in{\rm NTIME}[n^{k}]\subseteq\bigcup_{i\in\mathbb{N}_{1}}{\rm NTIME}[n^{i}].$
It is time for us to show that Definition A.3 is equivalent to Definition A.1. For the purposes of brevity, we omit the proof and refer the reader to [Kar72] (cf. Theorem 1 in [Kar72]), or to the proof of Theorem $7$.20202020 in [Sip13] which is a theorem answering the same question discussed here (see p. $294$ of [Sip13]). Thus this completes the proof.